### v3

# Aditya Tanna, Pratinav Seth, Mohamed Bouadi Utsav Avaiya, Vinay Kumar Sankarapu Lexsi Labs, India & France

Aditya Tanna, Pratinav Seth, Mohamed Bouadi
Utsav Avaiya, Vinay Kumar Sankarapu
Lexsi Labs, India & France
1  Introduction
Tabular data i.e. structured collections of records (rows) and attributes (columns)—underpins a vast array of real-world applications, from healthcare and finance to industrial analytics and scientific research. These datasets form the backbone of modern decision-making systems: hospitals rely on patient records for diagnostic risk prediction, financial institutions assess creditworthiness and fraud using structured features, and scientific studies aggregate multi-sensor measurements to discover new insights. Yet despite this ubiquity, deep learning on tabular data has long lagged behind the breakthroughs achieved in domains such as natural language processing and computer vision.
Traditional approaches like gradient-boosted decision trees (e.g., XGBoost [chen2016xgboost], LightGBM [ke2017lightgbm], CatBoost [prokhorenkova2018catboost]) remain dominant because they reliably handle heterogeneous features, variable sample sizes, and small datasets—challenges that standard neural networks often struggle with. Deep models frequently underperform on tabular benchmarks due to the lack of spatial or temporal structure, making feature extraction less amenable to convolutional or sequential architectures. Moreover, tabular datasets often exhibit high heterogeneity—mixing numerical, categorical, and ordinal variables—and suffer from missing values and skewed distributions, which violate many of the inductive biases that make deep learning successful in other modalities. As a result, practitioners continue to favor tree ensembles and gradient boosting for their robustness, interpretability, and low data requirements.
Recently, the advent of tabular foundation models (TFMs) has begun to reshape this landscape. These models extend the “foundation model” paradigm—pretraining on large, diverse corpora followed by downstream adaptation—to structured data domains. Models such as TabPFN [tabpfn, TabPFN2, tabpfn25], TabICL [tabicl], ContextTAB [contexttab], OrionMSP [orionmsp2025], and OrionBiX [orionbix2025] leverage large-scale pre-training on synthetic or massive tabular datasets to provide generic, out-of-the-box learning capabilities. These developments mark a significant step toward “general-purpose” tabular learners that can adapt rapidly to new datasets with minimal supervision.
However, despite their promise, adopting these models in real-world practice remains hampered by several frictions:
  • Diverse preprocessing requirements: Each model demands its own data encoding, feature normalization, and missing-value strategy—requiring practitioners to build bespoke pipelines for each model family. For example, TabPFN expects numerically encoded categorical features consistent with its synthetic priors, whereas TabICL and Orion architectures rely on learned embeddings or set-transformer encoders for categorical attributes.
  • Fragmented APIs and training protocols: Some models operate purely in zero-shot inference mode (zero fine-tuning), others support supervised fine-tuning (SFT) or parameter-efficient fine-tuning (PEFT), specifically Low-Rank Adaptation (LoRA; see Section 2 for details). Managing these disparate workflows is laborious and error-prone, particularly when comparing multiple model families under consistent experimental settings.
  • Evaluation gaps in deployment-relevant metrics: While accuracy improvements are well documented, aspects such as calibration of probability estimates, fairness across subgroups, and resource-efficiency trade-offs are under-explored in a unified framework. This absence of standardization makes it difficult to assess whether performance gains translate to trustworthy or deployable behavior.
  • Model selection complexity: With multiple models and tuning strategies available, users face non-trivial questions: Which model best suits a small vs. large dataset? What is the memory/latency trade-off? How do calibration and fairness behave under different tuning regimes? In practice, this uncertainty often discourages the adoption of TFMs despite their potential benefits.
To bridge these gaps, we introduce TabTune, a unified, scikit-learn-compatible toolkit that standardizes the full modelling workflow for tabular foundation models. TabTune offers:
  • A single interface for multiple TFM families, handling model-specific preprocessing internally and exposing consistent fit–predict semantics.
  • Support for a spectrum of adaptation strategies: zero-shot inference, meta-learning, full supervised fine-tuning (SFT), and parameter-efficient fine-tuning (PEFT). This design allows practitioners to switch between inference paradigms seamlessly using a unified API.
  • Built-in diagnostics for calibration (Expected Calibration Error, Maximum Calibration Error, Brier score) and fairness (Statistical Parity Difference, Equalised Odds Difference, Equalised Opportunity Difference), enabling a holistic assessment of model trustworthiness and risk.
  • A systematic benchmarking module across standard tabular suites (e.g., TALENT [talent], OpenML-CC18 [openmlcc18]) that evaluates accuracy, calibration, fairness, and resource efficiency in a unified ranking framework, facilitating reproducible and comparative studies of TFM performance.
With TabTune, practitioners can quickly experiment with multiple model/tuning configurations using consistent functions (e.g., .fit(), .predict(), .evaluate()), without rewriting preprocessing logic or training loops. Moreover, TabTune serves as an experimental bed for studying the interplay between learning paradigms—such as zero-shot inference versus parameter-efficient tuning—and their downstream effects on calibration, fairness, and computational efficiency. For detailed system architecture, see Section 2. Through comprehensive experiments, we demonstrate that TabTune not only simplifies workflows, but also provides new insights into the trade-off space of accuracy, calibration, fairness and efficiency across tabular foundation models.
In the remainder of the paper, Section 2 reviews prior work on tabular foundation models and adaptation techniques. Section 2 details the system design of TabTune, Section 6 describes our experimental setup and benchmark results, and Section 7 offers practical guidance on model selection, discusses limitations, and outlines future directions.
2  Related Work
2.1  Traditional Machine Learning and Deep Learning for Tabular Data
Tabular data—structured as records (rows) and features (columns)—is pervasive across applications such as healthcare, finance, industrial analytics, and scientific research. For decades, models based on gradient-boosted decision trees (GBDTs) such as XGBoost, LightGBM, and CatBoost have dominated supervised learning on tabular tasks. These methods reliably handle heterogeneous feature types, missing values, and small to medium sample sizes, while requiring only modest hyper-parameter tuning. Their robustness and simplicity have made them the de facto standard for tabular prediction tasks in both research and production settings.
In recent years, deep neural networks—such as feed-forward multilayer perceptrons, attention-based architectures, and hybrid tree–neural models—have been applied to tabular data [borisov2022tabular, somvanshi2024tabular]. Despite significant progress, these models have often struggled to consistently outperform GBDTs in real-world benchmarks. This gap arises from several challenges: the heterogeneity of tabular data, limited dataset sizes, complex feature–target relationships, and the absence of large-scale pre-training strategies analogous to those that revolutionized vision and language modeling.
2.2  Tabular Foundation Models
A major paradigm shift has recently emerged with the advent of tabular foundation models (TFMs)—large pretrained architectures that aim to generalize across diverse tasks, often through in-context learning (ICL) or single forward-pass inference without task-specific gradient updates. In the context of TFMs, zero-shot inference refers to the "fit then predict" paradigm: during the fit phase, the model receives training samples as context (without gradient-based parameter updates), and during the predict phase, it performs in-context learning through forward passes alone to make predictions on test samples. Unlike traditional machine learning where "fit" implies gradient updates and training, in TFMs the fit step simply provides context samples to the pretrained model, enabling rapid adaptation without requiring training time.
TabPFN (Tabular Prior-data Fitted Network) represents one of the earliest and most influential TFMs. It employs a transformer architecture trained on vast collections of synthetic tabular datasets, effectively learning to approximate Bayesian inference across varying generative priors. This pre-training strategy enables TabPFN to achieve strong few-shot and zero-shot performance on small to medium datasets (typically up to $\sim$10,000 samples), offering competitive accuracy without explicit fine-tuning [tabpfn, TabPFN2, tabpfn25]. Building on this foundation, TabICL extends the in-context learning paradigm to larger-scale tabular datasets by introducing a novel two-stage column-then-row attention mechanism. Through large-scale synthetic pre-training (up to $\sim$60K samples) and robust scaling to half a million samples at inference time, TabICL demonstrates that transformer-based ICL can challenge traditional tree ensembles even on large tabular domains [tabicl]. Another line of work, TabDPT, introduces an in-context learning (ICL)-based tabular foundation model trained with denoising-style, self-supervised objectives on real-world tables [tabdpt]. In parallel, other approaches explore the role of mixed synthetic priors, showing that combining causal, statistical, and randomized priors during pre-training can substantially improve downstream generalization and calibration performance across classification and regression tasks.
Collectively, these developments illustrate that, with sufficient architectural bias and large-scale pre-training, transformer-based TFMs can achieve broad generalization on structured data—challenging long-held assumptions about the superiority of tree-based ensembles in tabular learning. Recent comprehensive surveys have catalogued the landscape of tabular foundation models [jiang2025representation, badaro2023transformers], highlighting their potential to transform tabular learning. Recent work has demonstrated that careful fine-tuning strategies can yield significant improvements for TFMs [rubachev2025finetuning]. Meanwhile, researchers have argued that tabular foundation models should be a research priority [breugel2024position], given their potential impact across domains.
OrionMSP [orionmsp2025] introduces a transformer architecture for tabular in-context learning that combines multi-scale sparse attention—capturing both local and global feature dependencies efficiently—with a Perceiver-style latent memory that enables safe bidirectional communication between training and test samples. It uses column-wise embeddings via Set Transformers to model feature distributions and a split-masked causal attention mechanism for proper ICL reasoning. The design achieves near-linear attention scaling while maintaining high accuracy across diverse datasets, rivaling or outperforming TabICL and TabPFN—especially on heterogeneous, high-dimensional, and imbalanced datasets—thus demonstrating robust, efficient, and generalizable tabular learning without gradient updates. Similarly, OrionBiX [orionbix2025] employs a biaxial attention mechanism for tabular in-context learning, enabling bidirectional context modeling across both rows and columns of tabular data. This dual-axis attention pattern allows the model to capture complex feature interactions and sample relationships simultaneously, providing a more comprehensive representation of tabular structure. The biaxial design facilitates effective information flow between training and test samples, enabling strong in-context learning performance across diverse tabular datasets while maintaining computational efficiency.
2.3  Representation Learning and Alternative Approaches
While transformer-based TFMs dominate recent work, alternative representation learning paradigms have been explored. Graph neural networks have been adapted for tabular data, modeling feature relationships through graph structures [li2024gnn]. Extensions of foundation models to time series forecasting demonstrate their versatility; for instance, TabPFN-v2 has shown competitive performance on temporal tasks [hoo2025tables], suggesting that tabular foundation models may generalize beyond traditional classification settings.
2.4  Tabular ML Toolkits and Frameworks
Several frameworks have emerged to facilitate deep learning with tabular data, each addressing different aspects of the modeling pipeline. AutoGluon [erickson2020autogluon] is an AutoML framework that automates model selection and hyperparameter tuning for tabular data, employing neural architecture search and ensemble methods to achieve strong predictive performance with minimal manual configuration. PyTorch Tabular [joseph2021pytorch] provides a modular framework for deep learning with tabular data, offering reusable components for neural network architectures tailored to structured data. However, these frameworks primarily focus on training models from scratch or performing architecture search, rather than working with pretrained tabular foundation models.
In contrast, TabTune is the first unified framework specifically designed for tabular foundation models, bridging the gap between pretrained TFM architectures and practical deployment workflows. While existing toolkits excel at discovering or building models from scratch, TabTune addresses the unique challenges of working with pretrained TFMs: standardizing diverse preprocessing requirements, unifying adaptation strategies (zero-shot inference, meta-learning, supervised fine-tuning, and parameter-efficient fine-tuning), and providing built-in evaluation diagnostics for calibration and fairness. This specialization enables practitioners to leverage the powerful generalization capabilities of pretrained TFMs without the complexity of managing model-specific interfaces and workflows.
2.5  Adaptation Techniques: Meta-Learning, Fine-Tuning, and PEFT
Although inference-only TFMs exhibit strong generalization, many real-world applications benefit from adapting to target distributions. Three key adaptation paradigms have emerged.
Meta-learning (or episodic fine-tuning) uses support–query splits from training data to retain in-context generalization while improving task-specific performance [finn2017model]. This approach has proven effective for tabular foundation models such as TabICL, where episodic training enhances accuracy and supports few-shot adaptation to new distributions [tabicl, tabdpt].
Supervised fine-tuning (SFT) updates all model parameters on labeled data when sufficient resources are available [howard2018universal]. In tabular domains, full fine-tuning often yields notable gains over inference-only use, especially on large datasets where models learn domain-specific patterns [tabicl, tabdpt]. However, it remains computationally and storage intensive, as every parameter must be updated per task.
Parameter-efficient fine-tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA) [hu2021lora], limit updates to low-dimensional subspaces, greatly reducing compute and memory costs while preserving performance. Surveys categorize PEFT into additive (adapters), selective (partial updates), and reparameterization-based (low-rank) approaches [liu2024peft]. Empirical studies show that PEFT matches or surpasses full fine-tuning with far lower resource demands [razuvayevskaya2024peft]. For tabular foundation models, PEFT enables efficient domain adaptation under tight resource budgets, supporting rapid experimentation and scalable deployment. Together, these paradigms bridge the gap between zero-shot inference and fully supervised modeling, promoting adaptable TFMs for diverse applications.
2.6  Calibration, Fairness, and Responsible Deployment
As tabular foundation models progress from research to deployment, challenges of trustworthiness, calibration, fairness, and robustness have become increasingly central.
Calibration reliability—the alignment between predicted probabilities and actual outcomes—is vital for high-stakes decision-making. Guo et al. [guo2017calibration] showed that modern neural networks often exhibit systematic miscalibration, motivating metrics such as Expected Calibration Error (ECE) and Maximum Calibration Error (MCE) to measure these gaps. For tabular foundation models, early results suggest that pretrained transformers yield better-calibrated confidence estimates than classical ensembles, especially in low-data settings where in-context learning supports more nuanced uncertainty estimation.
Fairness requires that models avoid systematic bias across demographic groups [barocas2019fairness]. Fairness frameworks propose criteria such as equalized odds [hardt2016equality]—requiring equal true and false positive rates—and equal opportunity, focusing on true positive rate parity. [pleiss2017fairness] revealed a fundamental tension between perfect calibration and strict fairness, showing that these objectives can conflict. In-context learning setups may further amplify biases present in demonstration data. Mitigation strategies include group-balanced sampling, fair demonstration selection, and uncertainty-based data curation.
Robustness—the stability of model behavior under distributional shift, perturbations, and adversarial attacks—remains an open challenge. While tabular foundation models benefit from large-scale, heterogeneous pretraining, their behavior under shift and noise is not yet fully characterized. These considerations highlight the need for unified evaluation frameworks that jointly assess predictive performance, calibration, fairness, and robustness to ensure the responsible deployment of tabular foundation models.
3  System Design
Figure 1: Overview of the TabTune architecture. Raw tabular data are processed through a model-aware pre-processing engine and modular tuning pipeline, coordinated by a central orchestration engine with integrated evaluation assessment modules.
The design of TabTune emphasizes abstraction of the operational complexity associated with modern tabular foundation models (TFMs), while preserving flexibility for advanced adaptation, benchmarking, and experimentation. The framework adopts a strictly modular architecture in which preprocessing, tuning, and orchestration layers are decoupled yet interoperable, ensuring extensibility and transparency across research and production workflows (Figure 2).
3.1  Design Principles
Unified Interface. TabTune exposes a consistent, high-level API that provides unified access to all supported TFMs, including TabPFN, TabDPT, TabICL, Mitra, ContextTab, OrionMSP and OrionBiX . Through a shared syntax for model initialization, training, and evaluation, it eliminates the need for users to adapt to model-specific conventions or data representations, enabling seamless experimentation across architectures.
Model-Aware Automation. The framework automatically identifies the chosen model and configures preprocessing, hyperparameters, and adaptation procedures accordingly. Whether deploying a zero-shot TabPFN, performing episodic meta-learning with TabICL, or applying parameter-efficient LoRA fine-tuning, TabTune dynamically selects optimized defaults aligned with the model’s internal design and capacity. This automation reduces manual configuration overhead and ensures consistency across runs.
Unified Workflow Compatibility. All models in TabTune follow a consistent workflow based on standardized method calls such as fit(), predict(), and evaluate(). This design allows easy integration into broader machine learning pipelines and benchmarking frameworks, enabling reproducible and interpretable experimentation without requiring specialized wrappers or retraining scripts.
3.2  Modular Architecture
Figure 2: Modular architecture of TabTune. The TabularPipeline orchestrates the workflow by chaining the model-aware DataProcessor, the encapsulated TFM, and the adaptive TuningManager, forming a cohesive and control framework.
TabTune’s architecture is composed of four interdependent modules that collectively handle data processing, model adaptation, inference, and evaluation. These components are designed to remain logically independent yet harmonized through a unified orchestration layer.
  • TabularPipeline module, which serves as the main user interface and orchestrator of the modeling workflow. It encapsulates the selected TFM, manages persistence through save() and load() utilities, and coordinates the complete pipeline—from preprocessing and fine-tuning to evaluation and model comparison.
  • The DataProcessor component is responsible for handling model-specific preprocessing logic. Upon initialization, it dynamically loads specialized processors and performs imputation, normalization, encoding, and embedding generation as required.
  • The TuningManager module constitutes the computational backbone of TabTune. It executes the selected adaptation strategy—ranging from zero-shot inference to supervised fine-tuning, meta-learning, and parameter-efficient fine-tuning (PEFT). The manager implements training routines in a unified format, abstracts gradient updates, and integrates lightweight adaptation methods such as LoRA to reduce computational and memory costs while maintaining model generalization. For meta-learning configurations, episodic training is supported through dynamic support–query sampling, preserving the in-context adaptability of transformer-based models.
  • Complementing these components, the TabularLeaderboard module enables systematic evaluation and large-scale comparison of model variants. It automates training and evaluation over consistent data partitions and produces ranked summaries across metrics such as AUC, F1-score, and inference efficiency. This functionality transforms comparative model assessment into a reproducible and configurable experiment suite.
3.3  Supported Models
TabTune includes built-in support for seven prominent tabular foundation models, default tuning configurations, and aligned preprocessing pipelines. This design ensures that every model can be instantiated, fine-tuned, and evaluated with minimal configuration effort, while preserving model-specific fidelity.
Table 1: Tabular foundation models supported in TabTune. Paradigm indicates the learning approach—ICL, Prior-Fitted Network, Denoising Transformer, or Scalable ICL. Key Feature summarizes each model’s core architectural element (e.g., attention).
Model Paradigm Key Feature
OrionMSP In-Context Learning Multi-scale sparse attention with Perceiver-style latent memory
OrionBiX In-Context Learning Biaxial attention with bidirectional context modeling
TabICL In-Context Learning Two-stage column-then-row attention for large-scale tabular ICL
Mitra Scalable ICL 2D attention (row and column) with mixed synthetic priors
ContextTab Semantics-Aware ICL Utilizes modality-specific semantic embeddings
TabDPT In-Context Learning Employs diffusion-based self-supervised pre-training
TabPFN Prior-Fitted Network Approximates Bayesian inference on synthetic priors
4  Implementation
This section describes how TabTune’s implementation is organized around three core components—the DataProcessor, TuningManager, and TabularPipeline—which together enable seamless preprocessing, adaptation, and inference through a standardized interface.
4.1  Model-Aware Preprocessing
The DataProcessor dynamically loads model-specific routines to format data correctly with minimal user setup. For TabPFN, features are mapped to numerical or integer-encoded tensors to align with its synthetic priors. ContextTab [contexttab] adds semantic vectors for column names and categorical values. TabICL and Mitra apply normalization and model-tailored transformations. For OrionMSP [orionmsp2025] and OrionBiX [orionbix2025], preprocessing is distribution- and hierarchy-aware: columns are embedded with Set Transformer-based induced set attention on train rows only, numerical features are standardized and categorical values embedded, multi-scale row inputs and block-sparse attention masks are constructed, Perceiver-style memory is written by train rows and read by all rows, and label injection uses split-masked attention to prevent test-to-train leakage. Compared to TabICL [tabicl], ORION adds distribution-aware column embeddings and multi-scale sparsity for scalable, ICL-safe processing. Together, these routines ensure each TFM receives inputs in its expected schema.
4.2  Flexible Tuning Strategies
The TuningManager module governs model adaptation and optimization, executing the selected strategy specified during pipeline initialization. It provides a unified training controller that supports four complementary regimes—inference, supervised fine-tuning, meta-learning, and parameter-efficient fine-tuning (PEFT)—under a common abstraction.
  • Zero-shot inference, TabTune performs zero-shot prediction using pretrained weights, exploiting the generalization capacity of models such as TabPFN for small datasets without parameter updates.
  • Supervised fine-tuning extends this by optimizing all model parameters through gradient-based learning on labeled data, allowing users to specify training details (e.g., number of epochs or learning rate) via the configuration dictionary. For in-context learning models such as TabICL and Mitra, batches are split into pseudo-episodes where the first half serves as support context and the second half as query targets, maintaining the episodic structure required by the model architecture.
  • Meta-learning fine-tuning introduces episodic adaptation, where models are trained on dynamically sampled support–query pairs to maintain in-context generalization while improving task-specific accuracy. Each episode randomly samples disjoint support and query sets from the training data, with class labels normalized to contiguous indices within each episode to ensure compatibility with fixed-size classification heads. This approach is particularly effective for transformer-based architectures like TabICL and TabDPT that rely on in-context reasoning.
  • Parameter-Efficient fine-tuning (PEFT)—implemented via Low-Rank Adaptation (LoRA)—reduces computational cost by updating only small low-rank matrices within attention layers. LoRA is configured with rank $r=8$ and scaling factor $\alpha=16$, injecting trainable adapters into attention projection layers. When LoRA application fails due to architectural constraints, the framework automatically falls back to standard fine-tuning. This technique significantly lowers memory requirements while preserving performance and is supported for both supervised and meta-learning modes.
4.2.1  PEFT Implementation Details:
The fine-tuning framework determines target layers for parameter-efficient adaptation automatically based on model architecture. Updates are typically restricted to attention projections and core transformer components, while dynamic or task-specific heads are excluded. If no predefined targets are detected, adapters are applied to all eligible linear layers. When adapter injection is incompatible with the underlying architecture, the system reverts to standard full-parameter fine-tuning.
4.2.2  Episodic Training Mechanics:
For meta-learning adaptation, episodes are formed by sampling disjoint support and query sets from the training data. Each episode contains $S$ labeled support examples and $Q$ query examples used for evaluation. The model is conditioned on the support labels and optimized to predict the query labels within each episode. To ensure compatibility with fixed-size classification heads, class labels are remapped to contiguous indices $\{0,1,\ldots,K-1\}$ according to the support set. Episodes in which query samples include unseen classes are excluded to preserve consistency.
Table 2: Fine-tuning strategy support matrix across tabular foundation models. $\checkmark$ = full support; $\ast$ = experimental (may revert to base fine-tuning); – = not supported. SFT = Supervised Fine-Tuning; Meta = Meta-Learning; PEFT SFT = Parameter-Efficient Supervised Fine-Tuning; PEFT Meta = Parameter-Efficient Meta-Learning.
Model SFT Meta-Learning PEFT SFT PEFT Meta-Learning
OrionMSP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OrionBiX $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TabPFN $\checkmark$ $\checkmark$ $\ast$ $\ast$
TabICL $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
TabDPT $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Mitra $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ContextTab $\checkmark$ $\ast$
4.2.3  Data Sampling Strategies
The framework incorporates optional resampling techniques to address label imbalance and promote a well-diversified dataset for model context, while remaining independent of the core modeling workflow. Resampling is applied exclusively to the training split to avoid any data leakage into the validation or test sets.
Supported methods : The available options include: smote (Synthetic Minority Over-sampling Technique), random_over (random oversampling), random_under (random undersampling), tomek (Tomek links removal), kmeans (cluster centroids undersampling), and knn (neighborhood cleaning rule). These methods enable flexible control over class balance and support reproducible experimentation across diverse tabular datasets.
4.3  Evaluation Utilities
TabTune provides comprehensive evaluation capabilities across performance, calibration, and fairness. Beyond these standard metrics, TabTune also includes built-in utilities for assessing model calibration and fairness—critical dimensions for responsible deployment in high-stakes applications.
  • The evaluate() method computes standard classification metrics including accuracy, precision, recall, F1-score, and AUC-ROC, enabling rapid assessment of predictive performance across models and tuning strategies (Listing LABEL:lst:performance_eval).
  • The evaluate_calibration() method computes Expected Calibration Error (ECE), Maximum Calibration Error (MCE), and Brier score to quantify the reliability of predicted probabilities. These metrics reveal whether a model’s confidence estimates align with actual prediction accuracy, essential for decision-making systems in healthcare, finance, and autonomous applications.
  • The evaluate_fairness() method assesses demographic parity and equalized odds across sensitive attributes, measuring whether predictions exhibit systematic bias across subgroups. Users specify sensitive features (e.g., gender, race, age) and receive Statistical Parity Difference (SPD), Equalised Odds Difference (EOD), and more, enabling practitioners to identify and mitigate algorithmic bias before deployment.
4.4  Installation
TabTune is designed to be easily integrated into existing Python environments. The library is open-source and available on the Python Package Index (PyPI), ensuring straightforward installation for practitioners.
To install the latest stable release, users can execute the following command:
1pip install tabtune
Alternatively, for the latest development version, the package can be installed directly from the source repository:
1git clone https://github.com/Lexsi-Labs/TabTune.git
2cd TabTune
3pip install -r requirements.txt
4pip install -e .
4.5  Pipeline Abstraction and Workflow
The TabularPipeline serves as the primary orchestration layer for integrating the preprocessing and tuning components. It manages the end-to-end flow—data ingestion, preprocessing, adaptation, and prediction—through a unified API that abstracts low-level operations.
The following demonstrates the core workflow for model initialization, training, and inference. The pipeline detects the model type, loads the corresponding preprocessing backend, and delegates adaptation to the TuningManager. In this example, the user specifies only the model and adaptation mode. TabTune internally triggers three automated steps: (1) the DataProcessor formats the dataset to match the model’s expected schema, (2) the TuningManager selects and executes the correct adaptation strategy, and (3) the TabularPipeline aggregates predictions and handles evaluation. This design illustrates how the framework abstracts operational complexity behind a minimal interface, enabling quick prototyping and reproducible experimentation.
1from tabtune import TabularPipeline
2# Initialize the pipeline with model configuration
3pipeline = TabularPipeline(model_name="OrionMSP",tuning_strategy="inference")
4# Train and predict
5pipeline.fit(X_train, y_train)
6predictions = pipeline.predict(X_test)
4.6  Benchmarking with TabularLeaderboard
To facilitate large-scale model comparison and reproducibility, TabTune includes the TabularLeaderboard module. This component automates benchmarking across multiple TFMs and adaptation modes under consistent data splits, producing ranked summaries based on selected metrics such as AUC, F1-score, and training efficiency. Listing LABEL:lst:leaderboard shows a representative use case where several models and tuning regimes are benchmarked on the same dataset. Each configuration defines its own adaptation strategy and hyperparameters, while TabTune ensures standardized preprocessing and evaluation across all runs.The leaderboard system illustrates the high-level orchestration capabilities of TabTune’s architecture. Internally, it constructs multiple TabularPipeline instances, each managing its respective preprocessing and tuning configuration, and executes them in parallel under uniform conditions. The results are aggregated and ranked automatically, providing a reproducible, interpretable, and scalable benchmarking framework for evaluating tabular foundation models. Following demonstrates both evaluation modes, which can be applied to any TabTune model after training.
1from tabtune import TabularLeaderboard
2# Initialize leaderboard
3leaderboard = TabularLeaderboard(X_train, X_test, y_train, y_test)
4# Add model configurations
5leaderboard.add_model(model_name=’TabPFN’,tuning_strategy=’inference’)
6leaderboard.add_model(model_name=’OrionBix’,tuning_strategy=’finetune’,
7 tuning_params={’epochs’: 5})
8leaderboard.add_model(model_name=’OrionMSP’,tuning_strategy=’inference’)
9# Execute and rank models
10results = leaderboard.run(rank_by=’roc_auc_score’)
5  Usage Scenarios
This section demonstrates TabTune’s capabilities through several scenarios, showcasing the framework’s functionality. These examples provide hands-on illustrations of the system components described in Section 4.
5.1  Basic Usage and Multi-Model Switching
TabTune’s unified API allows seamless experimentation across foundation models with identical workflows. Switching between the seven supported models requires only a parameter change—no need to learn new APIs or preprocessing steps. This consistency enables rapid prototyping, performance comparison, and deployment without rewriting data or training code. Model persistence ensures reproducibility and version control.
1from tabtune import TabularPipeline
2# Example 1: TabPFN Model
3pipeline_pfn = TabularPipeline(model_name="TabPFN",tuning_strategy="inference")
4pipeline_pfn.fit(X_train, y_train)
5predictions_pfn = pipeline_pfn.predict(X_test)
6# Example 2: OrionMSP Model with semantic features
7pipeline_OrionMSP = TabularPipeline(model_name="OrionMSP",tuning_strategy="inference")
8pipeline_OrionMSP.fit(X_train, y_train)
9predictions_OrionMSP = pipeline_OrionMSP.predict(X_test)
5.2  Model Comparison and Results Analysis
This scenario showcases systematic benchmarking of multiple models and tuning strategies with detailed performance insights. The leaderboard automates cross-validation, metric computation, and ranking across all configurations. Results can be exported as structured DataFrames for analysis, visualization, or automated model selection—enabling data-driven decisions on the best model–tuning combination for a given dataset and deployment setting.
1from tabtune import TabularLeaderboard
2leaderboard = TabularLeaderboard(X_train, X_test,y_train, y_test)
3# Add multiple model configurations
4leaderboard.add_model(model_name=’TabPFN’,tuning_strategy=’inference’)
5leaderboard.add_model(model_name=’TabICL’,tuning_strategy=’finetune’,tuning_params={’epochs’:5})
6leaderboard.add_model(model_name=’OrionMSP’,tuning_strategy=’inference’)
7# Run comparison and get detailed results
8results = leaderboard.run(rank_by=’roc_auc_score’)
5.3  Fine-Tuning Demonstration
This scenario demonstrates the different fine-tuning strategies available in TabTune. Each approach offers distinct trade-offs between performance, and generalization capability.
5.3.1  Supervised Fine-Tuning (SFT)
Supervised fine-tuning optimizes all model parameters using standard mini-batch gradient descent. This approach typically achieves the highest accuracy when sufficient training data is available.
1from tabtune import TabularPipeline
2# Full supervised fine-tuning
3sft_pipeline = TabularPipeline(
4 model_name=’OrionMSP’,
5 tuning_strategy=’finetune’,
6 tuning_params={’finetune_mode’: ’sft’,’epochs’: 5,’learning_rate’: 1e-4}
7)
8sft_pipeline.fit(X_train, y_train)
9sft_metrics = sft_pipeline.evaluate(X_test, y_test)
5.3.2  Meta-Learning Fine-Tuning
Meta-learning fine-tuning uses episodic training, where the model learns to adapt from support sets to query sets within each episode. This approach preserves in-context learning capabilities while improving task-specific performance.
1from tabtune import TabularPipeline
2# Meta-learning fine-tuning
3meta_pipeline = TabularPipeline(
4 model_name=’OrionMSP’,
5 tuning_strategy=’finetune’,
6 tuning_params={
7 ’finetune_mode’: ’meta-learning’,
8 ’epochs’: 3, ’learning_rate’: 5e-5, ’support_size’: 48, ’query_size’: 32
9 }
10)
11meta_pipeline.fit(X_train, y_train)
12meta_metrics = meta_pipeline.evaluate(X_test, y_test)
5.3.3  Parameter-Efficient Fine-Tuning with SFT (PEFT SFT)
PEFT SFT combines LoRA adapters with supervised fine-tuning, achieving comparable performance to full fine-tuning while reducing memory usage by 60–80%.
1from tabtune import TabularPipeline
2# PEFT with supervised fine-tuning
3peft_sft_pipeline = TabularPipeline(
4 model_name=’TabICL’,
5 tuning_strategy=’peft’,
6 tuning_params={
7 ’finetune_mode’: ’sft’, ’epochs’: 5,
8 ’peft_config’: {’r’: 8,’lora_alpha’: 16,’lora_dropout’: 0.05}
9 }
10)
11peft_sft_pipeline.fit(X_train, y_train)
12peft_sft_metrics = peft_sft_pipeline.evaluate(X_test, y_test)
5.3.4  Parameter-Efficient Fine-Tuning with Meta-Learning (PEFT Meta-Learning)
PEFT meta-learning combines LoRA adapters with episodic training, offering the benefits of both parameter efficiency and in-context generalization.
1from tabtune import TabularPipeline
2# PEFT with meta-learning
3peft_meta_pipeline = TabularPipeline(
4 model_name=’TabICL’,
5 tuning_strategy=’peft’,
6 tuning_params={’finetune_mode’: ’meta-learning’, ’epochs’: 3,
7 ’support_size’: 48, ’query_size’: 32,
8 ’peft_config’: {’r’: 8, ’lora_alpha’: 16, ’lora_dropout’: 0.05}}
9)
10peft_meta_pipeline.fit(X_train, y_train)
11peft_meta_metrics = peft_meta_pipeline.evaluate(X_test, y_test)
5.4  Comprehensive Model Evaluation
This scenario demonstrates TabTune’s built-in evaluation utilities across three dimensions: standard performance metrics, probability calibration, and fairness analysis.
5.4.1  Performance Metrics Evaluation
Standard classification metrics provide the foundation for model assessment. TabTune’s evaluate() method computes accuracy, precision, recall, F1-score, and AUC-ROC in a unified call.
1from tabtune import TabularPipeline
2pipeline = TabularPipeline(model_name=’TabICL’,tuning_strategy=’finetune’)
3pipeline.fit(X_train, y_train)
4# Comprehensive performance evaluation
5performance = pipeline.evaluate(X_test, y_test)
6print(f"Accuracy: {performance[’accuracy’]:.4f}")
7print(f"AUC-ROC: {performance[’roc_auc_score’]:.4f}")
8print(f"F1-Score: {performance[’f1_score’]:.4f}")
5.4.2  Calibration Evaluation
Calibration assessment quantifies the reliability of predicted probabilities—essential for deployment in decision-critical applications. The evaluate_calibration() method returns ECE, MCE, and Brier score metrics.
1# Assess probability calibration quality
2calibration_metrics = pipeline.evaluate_calibration(X_test, y_test, n_bins=15)
3print(f"Expected Calibration Error: {calibration_metrics[’expected_calibration_error’]:.4f}")
4print(f"Maximum Calibration Error: {calibration_metrics[’maximum_calibration_error’]:.4f}")
5print(f"Brier Score: {calibration_metrics[’brier_score_loss’]:.4f}")
5.4.3  Fairness Evaluation
Fairness analysis detects systematic bias across demographic groups, enabling responsible model deployment. Users specify sensitive attributes, and TabTune computes demographic parity and equalized odds metrics.
1# Evaluate fairness across sensitive attributes
2fairness_metrics = pipeline.evaluate_fairness(X_test,y_test,sensitive_features=gender_column)
3print(f"Statistical Parity Difference: {fairness_metrics[’statistical_parity_difference’]:.4f}")
4print(f"Equalized Odds Difference: {fairness_metrics[’equalized_odds_difference’]:.4f}")
5.5  Model Persistence and Checkpointing
TabTune supports end-to-end pipeline persistence and checkpointing for reproducible experiments and deployment.
1from tabtune import TabularPipeline
2# Train and save a fine-tuned model
3pipeline = TabularPipeline( model_name=’TabICL’,tuning_strategy=’finetune’,tuning_params={’epochs’: 5}
4)
5pipeline.fit(X_train, y_train)
6performance = pipeline.evaluate(X_test, y_test)
7print(performance)
8# Save the entire pipeline (preprocessor + model + config)
9pipeline.save(’my_model.joblib’)
10# Load and use for predictions
11loaded_pipeline = TabularPipeline.load(’my_model.joblib’)
12performance = loaded_pipeline.evaluate(X_test, y_test)
13print(performance)
6  Experimental Results
We conduct evaluation of TabTune across multiple dimensions: predictive performance, probability calibration, fairness, and scalability. This section presents our experimental design and detailed results across diverse scenarios.
6.1  Experimental Design
Benchmark Suites and Datasets : Our experiments encompass three well-established benchmark suites i.e. TALENT [talent] (181 datasets), OpenML-CC18 [openmlcc18] (72 datasets), and TabZilla [tabzilla] (36 datasets). Facilitating a systematic comparison across a broad spectrum of tabular learning settings. We further conduct domain-specific evaluations in high-impact areas such as medical and financial domains to assess the practical applicability of the methods. All experiments adhere to the standardized dataset splits defined by each benchmark, ensuring both reproducibility and fairness in comparison. To maintain consistency and fairness across model families, benchmark results are computed only on the common subset of datasets available for all evaluated models within each benchmark suite. This unified evaluation protocol ensures that performance rankings and metrics reflect genuine methodological differences rather than disparities in dataset coverage. In total, our evaluations span 155/181 datasets for TALENT, 27/36 datasets for TabZilla, and 63/72 datasets for OpenML-CC18. Some datasets were excluded due to out-of-memory (OOM) errors even on H200 GPUs or CUDA-related issues, particularly affecting TabPFN-based models.
Models and Adaptation Strategies: We evaluate recent tabular foundation models—TabPFN, TabICL, OrionMSP, OrionBiX, Mitra, ContextTab, and TabDPT—under multiple adaptation paradigms, including zero-shot inference, meta-learning fine-tuning, supervised fine-tuning, parameter-efficient supervised fine-tuning, and parameter-efficient meta-learning. In addition, we include established traditional baselines using AutoGluon [erickson2020autogluon] such as XGBoost, LightGBM, CatBoost, and Random Forest as strong reference models for comparison.
Evaluation Metrics : Our assessment covers four complementary dimensions:
Performance: Classification Accuracy (ACC), AUC-ROC, and weighted F1-score (F1) quantify predictive capability across standard benchmark datasets (TALENT, OpenML-CC18, TabZilla) covering diverse dataset characteristics including small and large sample sizes, narrow and wide feature spaces, and balanced and imbalanced class distributions. It is important to clarify how mean rank values are derived. Within each benchmark suite, models are ranked by accuracy on every dataset (lower rank = better performance), and these per-dataset ranks are averaged to obtain the overall mean rank. Thus, a lower mean rank indicates stronger and more consistent performance across datasets, rather than the highest score on any single task. While absolute metrics (accuracy, F1) reflect peak task-level performance, mean rank provides a normalized measure of cross-dataset generalization consistency.
Scalability: Analysis of performance variation across dataset sizes, feature dimensionality, and class imbalance provides practical guidance for model selection. This analysis uses the same benchmark datasets as performance evaluation, aggregated and summarized across these dimensions to reveal scalability patterns.
Calibration: We evaluate how fine-tuning impacts calibration metrics across different models and adaptation strategies, measuring the alignment between predicted probabilities and actual outcomes, which is critical for decision-making systems. Our evaluation uses three metrics: (I) Expected Calibration Error (ECE), which measures the average deviation between predicted confidence and observed accuracy across binned predictions; (II) Maximum Calibration Error (MCE), which captures the worst-case calibration gap, identifying regions of systematic miscalibration; and (III) Brier score, which quantifies both calibration and refinement by penalizing squared deviations between predicted probabilities and binary outcomes.
Fairness: We assess equitable treatment across demographic subgroups using established fairness-oriented datasets with sensitive attributes (gender, race). We measure three fairness criteria that capture different notions of equitable treatment: (I) Statistical Parity Difference (SPD) quantifies the absolute difference in positive prediction rates across groups, testing whether the model provides equal opportunity independent of sensitive attributes. A model satisfies demographic parity when SPD approaches zero, indicating that positive predictions are distributed proportionally across demographic groups. (II) Equalized Odds Difference (EOD) measures disparities in both true positive and false positive rates across groups, ensuring consistent error profiles across subgroups. This criterion is more stringent than statistical parity, requiring not only proportional positive predictions but also equal accuracy within each group. (III) Equalized Opportunity Difference (EOpD) focuses specifically on true positive rate parity, ensuring that qualified individuals receive equal treatment regardless of group membership. This metric is particularly relevant in settings where false negatives have severe consequences, such as loan approvals or hiring decisions.
Analysis Focus : A central goal of our evaluation is to analyze how different adaptation strategies—zero-shot inference, meta-learning fine-tuning, and supervised fine-tuning—affect model behavior across multiple dimensions. We examine whether fine-tuning improves performance, influences calibration, and impacts fairness across domains. This comparative analysis offers practical guidance on selecting appropriate adaptation strategies for tabular foundation models in different application contexts. Detailed dataset statistics and information refer to Appendix 10.
All Experiments are executed on NVIDIA L40S (48Gb VRAM), with H200 (141Gb VRAM) used for memory-intensive cases. This ensures consistent execution across all experiments while handling the computational demands of models.
6.2  Performance Evaluation
Table 3: Overall leaderboard across three benchmark suites—TALENT, OpenML-CC18, and TabZilla. Ranks denote the mean rank per benchmark suite (lower is better). Metrics: ACC = Accuracy, F1 = Weighted F1. The “All” column reports the aggregated rank across all suites for a strategy. Models are grouped by adaptation strategy. Formatting: 1st place; 2nd place within each group.
Models / Strategy All TALENT OpenML-CC18 TabZilla
Rank Rank ACC F1 Rank ACC F1 Rank ACC F1
Baselines + Inference (Zero-Shot)
XGBoost 7.94 7.43 0.8395 0.8351 6.57 0.8558 0.8537 6.82 0.8612 0.8326
CatBoost 7.69 6.90 0.8329 0.8252 6.98 0.8588 0.8520 6.70 0.8579 0.8384
Random Forest 8.71 7.56 0.8278 0.8201 7.10 0.8547 0.8497 9.27 0.8358 0.8399
LightGBM 8.03 7.51 0.8324 0.8238 6.89 0.8581 0.8493 5.80 0.8618 0.8211
TabICL 5.92 5.09 0.8465 0.8372 5.30 0.8667 0.8623 6.65 0.8734 0.8698
OrionBix 6.52 5.87 0.8339 0.8253 5.63 0.8653 0.8596 5.47 0.8728 0.8628
OrionMSP 4.61 4.06 0.8455 0.8353 4.58 0.8722 0.8676 4.36 0.8821 0.8786
TabPFN-2.0 5.45 4.48 0.8508 0.8406 5.48 0.8714 0.8663 5.65 0.8752 0.8716
TabPFN-2.5 4.13 3.62 0.8540 0.8440 4.36 0.8726 0.8672 3.97 0.8824 0.8791
Mitra 13.07 11.63 0.3935 0.2884 11.50 0.3614 0.2522 12.21 0.3152 0.1830
ContextTab 7.75 6.52 0.8373 0.8269 7.01 0.8639 0.8581 8.02 0.8389 0.8334
TabDPT 6.45 6.33 0.8402 0.8311 5.18 0.8672 0.8625 4.57 0.8814 0.8775
FineTune – Meta Learning
TabICL 3.66 3.33 0.8344 0.8253 3.05 0.8664 0.8597 4.15 0.6956 0.6845
OrionBiX 3.96 4.25 0.8158 0.8060 2.86 0.8548 0.8516 2.00 0.8726 0.8662
OrionMSP 2.26 1.80 0.8401 0.8310 2.82 0.8548 0.8516 1.73 0.8735 0.8672
TabPFN-2.0 2.42 2.07 0.8517 0.8414 2.42 0.8842 0.8784 2.42 0.8663 0.8603
Mitra 6.09 5.79 0.6416 0.5874 5.42 0.6164 0.5651 4.70 0.5592 0.5147
TabDPT 3.95 3.72 0.8255 0.8167 4.40 0.8534 0.8501 2.62 0.8500 0.8456
FineTune – Supervised Fine Tuning
TabICL 5.05 4.74 0.7668 0.7439 4.36 0.6838 0.6299 4.65 0.5670 0.4733
OrionBiX 4.21 3.54 0.7698 0.7469 4.30 0.7126 0.6595 4.25 0.6476 0.5782
OrionMSP 2.88 2.27 0.7908 0.7653 3.18 0.7995 0.7668 2.88 0.7454 0.7222
TabPFN-2.0 1.97 1.83 0.8459 0.8350 1.89 0.8697 0.8617 1.86 0.8433 0.8327
Mitra 5.98 5.62 0.5460 0.4382 5.02 0.5408 0.4309 5.52 0.4608 0.3467
TabDPT 2.79 2.98 0.8202 0.8094 2.25 0.8499 0.8424 1.81 0.8337 0.8260
FineTune – PEFT - Meta Learning
TabICL 4.32 4.07 0.7017 0.6867 4.07 0.7773 0.7605 3.45 0.7116 0.7003
OrionBiX 2.77 2.39 0.7854 0.7762 2.50 0.8471 0.8430 3.52 0.7370 0.7200
OrionMSP 2.21 2.25 0.7879 0.7728 1.93 0.8566 0.8453 1.33 0.8594 0.8581
Mitra 4.64 4.17 0.6369 0.5905 4.32 0.6000 0.5426 4.36 0.5461 0.4960
TabDPT 2.28 2.11 0.8002 0.7910 2.17 0.8600 0.8539 2.33 0.8495 0.8481
FineTune – PEFT - Supervised Fine Tuning
TabICL 3.70 3.72 0.5692 0.4638 3.06 0.8174 0.7965 2.92 0.7920 0.7776
OrionBiX 3.99 3.57 0.6358 0.5570 3.67 0.7380 0.6789 3.95 0.6707 0.6071
OrionMSP 2.01 2.01 0.6749 0.5956 1.73 0.8241 0.8033 1.47 0.8214 0.8071
Mitra 4.62 4.01 0.5294 0.4303 4.49 0.4965 0.3917 4.70 0.4606 0.3597
TabDPT 1.91 1.65 0.7988 0.7842 2.03 0.8500 0.8398 1.95 0.8461 0.8461
6.2.1  Overall Leaderboards
Table 3 presents comprehensive results across the TALENT, OpenML-CC18, and TabZilla benchmark suites, reporting mean rank, accuracy (ACC), and weighted F1-score (F1) for each model and adaptation strategy. The aggregated findings reveal clear and consistent trends in how adaptation mechanisms influence predictive performance and cross-dataset generalization.
Baselines and Zero-Shot Inference : Classical Machine Learning methods remain strong reference baselines, attaining mean accuracies between 0.833 and 0.861 with aggregated ranks around 6.0. In contrast, pretrained tabular foundation models (TFMs) exhibit markedly stronger generalization even without task-specific training.
Among pretrained TFMs, TabPFN-2.5 now achieves the best overall zero-shot rank (4.13), with accuracies of (ACC=0.8540/F1=0.8440) on TALENT, (ACC=0.8726/F1=0.8672) on OpenML-CC18, and (ACC=0.8824/F1=0.8791) on TabZilla. OrionMSP follows closely (overall rank 4.61), reaching (ACC=0.8455 / 0.8353) on TALENT, (ACC=0.8722/F1=0.8676) on OpenML-CC18, and (ACC=0.8821/F1=0.8786) on TabZilla. TabICL and TabDPT remain competitive (ranks 5.92 and 6.45, ACC $\approx$ 0.84–0.88), while Mitra and ContextTab lag behind. Overall, zero-shot TFMs exceed traditional machine-learning baselines by roughly 2–4 percentage points in accuracy and three to four rank positions, demonstrating the strength of pretrained priors.
Meta-Learning Fine-Tuning : Episodic meta-learning substantially improves both accuracy and rank stability relative to zero-shot inference. OrionMSP attains the lowest aggregate rank of 2.26, achieving (ACC=0.8401/F1=0.8310) on TALENT, 0.8548/0.8516 on OpenML-CC18, and (0.8735/0.8672) on TabZilla. TabPFN ranks 2.42 overall with accuracies of (0.8517/0.8414) on TALENT, (0.8842/0.8784) on OpenML-CC18, and (0.8663/0.8603) on TabZilla. TabDPT (rank 3.95) maintains consistent performance near (0.83/0.82). By contrast, architectures such as TabICL and OrionBiX exhibit instability on TabZilla, where TabICL’s F1 drops to 0.6845. These results indicate that meta-learning offers the most balanced compromise between generalization and task-specific adaptation.
Supervised Fine-Tuning (SFT) : Full parameter fine-tuning yields divergent outcomes across architectures. TabPFN benefits most, achieving the best overall rank (1.97) and maintaining high accuracy across all benchmarks (ACC = 0.8459–0.8697, F1 = 0.8350–0.8617). TabDPT ranks 2.79 overall and performs particularly well on TabZilla (rank 1.81), with (ACC = 0.8337, F1 = 0.8260). OrionMSP remains competitive (rank 2.88) with (ACC $\approx$ 0.79, F1 $\approx$ 0.76), whereas TabICL and OrionBiX suffer severe degradation—TabICL’s accuracy on TabZilla falls from 0.873 to 0.567 and F1 from 0.870 to 0.473 —indicating overfitting and loss of pretrained priors. These findings suggest that SFT is advantageous for Bayesian or probabilistic architectures such as TabPFN.
Parameter-Efficient Fine-Tuning (PEFT) : PEFT achieves near-SFT accuracy with significantly reduced computational overhead. In the meta-learning configuration, OrionMSP attains the best overall rank (2.21), recording (ACC = 0.7879/F1 = 0.7728) on TALENT, (0.8566/0.8453) on OpenML-CC18, and (0.8594/0.8581) on TabZilla. TabDPT follows closely (rank 2.28) with (0.8002/0.7910) on TALENT and (0.8600/0.8539) on OpenML-CC18. Under the supervised PEFT variant, TabDPT achieves the overall best rank (1.91) with (ACC $\approx$ 0.85 and F1 $\approx$ 0.84), while OrionMSP (rank 2.01) attains (0.821 / 0.807) on TabZilla. These results show that PEFT retains roughly 95% of full-fine-tuning accuracy while substantially lowering resource demands.
Discussion : Across all adaptation regimes, TFMs consistently outperform traditional baselines by 2 % points in accuracy and approximately four rank positions. TabPFN proves most resilient, achieving the top overall rank under SFT, whereas OrionMSP demonstrates superior cross-dataset generalization in meta-learning. TabDPT delivers near-optimal performance under PEFT, offering the most favorable efficiency–accuracy balance. Conversely, models such as TabICL and OrionBiX are highly sensitive to full parameter updates and prone to overfitting. Collectively, these observations establish that careful alignment between model architecture and adaptation strategy is critical for maximizing the performance of tabular foundation models.
Key Takeaways:
  • OVERALL LEADERBOARD: Across all benchmark suites, the strongest overall performers are TabPFN and OrionMSP. The TabPFN family (including TabPFN-2.0 and TabPFN-2.5) achieves the best aggregate results, with mean accuracies between 0.85–0.88 and weighted F1 scores of 0.84–0.88, reflecting exceptional cross-benchmark consistency. OrionMSP follows closely (overall rank 4.61 in zero-shot and 2.26 in meta-learning), reaching ACC $\approx 0.84$$0.88$ and F1 $\approx 0.83$$0.88$, with particularly strong generalization on OpenML-CC18 and TabZilla.
  • ZERO-SHOT PERFORMANCE: The TabPFN family and OrionMSP are the strongest zero-shot models. TabPFN-2.5 attains the best aggregate zero-shot rank (4.13), with ACC $\approx\textsc{0.85}$0.88 across TALENT, OpenML-CC18, and TabZilla. OrionMSP is a close second (rank 4.61), matching TabPFN’s accuracy on OpenML-CC18 and TabZilla while slightly trailing on TALENT. Both substantially outperform classical baselines by about 2–4 accuracy points.
  • SFT trade-offs: Full supervised fine-tuning benefits probabilistic architectures such as TabPFN, which achieves the highest accuracies ($\text{ACC}\approx 0.85$$0.87$, $\text{F1}\approx 0.83$$0.86$) across all benchmarks. In contrast, models such as TabICL and OrionBiX exhibit severe degradation under SFT, reflecting overfitting and loss of pretrained priors.
  • PEFT performance: Parameter-efficient fine-tuning attains near–full-finetuning accuracy. TabDPT achieves top results ($\text{ACC}\approx 0.85$, $\text{F1}\approx 0.84$) and maintains strong stability across benchmarks, while OrionMSP delivers competitive results ($\text{ACC}\approx 0.82$$0.86$, $\text{F1}\approx 0.80$$0.85$) with cross-dataset generalization.
6.2.2  Scalability Analysis
Table 4: Performance variation by dataset size across all benchmark suites. ACC = Accuracy; F1 = Weighted F1-score, averaged across datasets within each size category: Small (<<1K samples), Medium (1K–10K), and Large (>>10K). Values are on a 0–1 scale (higher is better). Models are grouped by adaptation strategy. Formatting: 1st place ; 2nd place within each group.
Models Small (<1K) Medium (1K-10K) Large (>10K)
Rank ACC F1 Rank ACC F1 Rank ACC F1
Baselines + Zero-Shot Inference
XGBoost 8.66 0.8168 0.7964 8.28 0.8363 0.8314 6.52 0.8961 0.8911
CatBoost 8.88 0.8124 0.7935 7.89 0.8340 0.8264 6.43 0.8786 0.8721
Random Forest 9.48 0.7988 0.8187 8.57 0.8285 0.8221 8.68 0.8680 0.8614
LightGBM 8.80 0.8143 0.7789 8.30 0.8314 0.8226 6.82 0.8817 0.8753
TabICL 6.96 0.8301 0.8338 5.86 0.8486 0.8398 5.42 0.8790 0.8731
OrionBiX 7.34 0.8330 0.8150 6.71 0.8348 0.8260 5.54 0.8716 0.8656
OrionMSP 6.76 0.8232 0.8194 4.53 0.8494 0.8402 3.66 0.8831 0.8754
TabPFN-2.0 7.42 0.8325 0.8131 4.63 0.8557 0.8462 6.75 0.8771 0.8701
TabPFN-2.5 5.78 0.8349 0.8194 4.07 0.8566 0.8470 3.42 0.8887 0.8831
Mitra 15.24 0.4334 0.3236 12.89 0.3600 0.2553 12.40 0.3879 0.2801
ContextTab 9.76 0.8179 0.8085 6.82 0.8430 0.8334 9.34 0.8548 0.8464
TabDPT 6.28 0.8333 0.8271 6.58 0.8424 0.8339 6.18 0.8821 0.8754
FineTune – Meta Learning
TabICL 4.68 0.7912 0.7806 3.64 0.8342 0.8256 3.06 0.8431 0.8360
OrionBiX 2.78 0.8149 0.8123 3.99 0.8283 0.8204 4.50 0.8410 0.8306
OrionMSP 3.21 0.8211 0.8151 2.34 0.8413 0.8330 1.38 0.8787 0.8721
TabPFN-2.0 3.31 0.8336 0.8256 2.13 0.8638 0.8548 2.76 0.8748 0.8668
Mitra 6.39 0.6417 0.5988 6.10 0.6083 0.5516 5.82 0.6585 0.6132
TabDPT 3.85 0.8224 0.8188 4.09 0.8289 0.8212 3.54 0.8682 0.8613
FineTune - Supervised FineTuning
TabICL 5.65 0.6746 0.6225 4.86 0.7553 0.7187 5.25 0.6693 0.6212
OrionBiX 5.20 0.7158 0.6792 4.04 0.7718 0.7386 4.03 0.6697 0.6272
OrionMSP 2.90 0.8135 0.7920 2.85 0.8036 0.7798 2.96 0.7231 0.6828
TabPFN-2.0 2.68 0.8216 0.8109 1.93 0.8580 0.8485 1.57 0.8572 0.8459
Mitra 6.48 0.5625 0.4408 6.05 0.5133 0.4116 5.42 0.5720 0.4501
TabDPT 2.66 0.8227 0.8153 2.87 0.8377 0.8284 2.65 0.8120 0.8002
FineTune – PEFT - Meta Learrning
TabICL 4.80 0.7518 0.7397 4.25 0.7251 0.7114 4.29 0.6962 0.6799
OrionBiX 3.33 0.8093 0.8042 2.86 0.7995 0.7939 2.29 0.7848 0.7665
OrionMSP 2.28 0.8275 0.8213 2.10 0.8134 0.8019 2.45 0.8013 0.7817
Mitra 5.21 0.6510 0.6012 4.57 0.6089 0.5599 4.55 0.6304 0.5795
TabDPT 2.92 0.8191 0.8157 2.38 0.8121 0.8049 1.83 0.8407 0.8299
FineTune - PEFT - Supervised FineTuning
TabICL 3.53 0.7957 0.7710 3.73 0.6424 0.5686 3.71 0.6084 0.5020
OrionBiX 4.38 0.7576 0.7102 3.99 0.6585 0.5892 3.75 0.6327 0.5381
OrionMSP 1.73 0.8298 0.8132 2.06 0.7190 0.6597 2.04 0.6931 0.6152
Mitra 5.53 0.5244 0.4133 4.59 0.5008 0.4013 3.98 0.5459 0.4476
TabDPT 2.85 0.8269 0.8182 1.67 0.8332 0.8245 2.02 0.7686 0.7418
We analyze how model performance varies with dataset size, feature dimensionality, and class imbalance to inform model selection under different data regimes to highlight which models and strategies perform best across varying data constraints.
Based on Dataset Size
Table 4 shows how fine-tuning influences performance across dataset scales.
  • On small datasets (<<1K samples), TabPFN-2.5 achieves the highest zero-shot accuracy (ACC=0.8349, F1=0.8194), marginally ahead of TabPFN-2.0 under meta-learning (0.8336/0.8256) and TabDPT (0.8333/0.8271). OrionMSP (PEFT–Meta, 0.8275/0.8213) follows closely, confirming that transformer backbones still generalize well under constrained data. Fine-tuning TabICL or OrionBiX on small datasets causes sharp degradation (ACC drops >>15%), revealing overfitting risk.
  • For medium-sized datasets (1K–10K), TabPFN in meta-learning configuration leads (ACC=0.8638, F1=0.8548), followed by its SFT variant (0.8580/0.8485); both outperform all classical baselines.
  • On large-scale datasets (>>10K), tree-based ensembles remain strong—XGBoost reaches (0.8911/0.8961)—but transformer TFMs scale competitively, with OrionMSP zero-shot (0.8754/0.8831) and TabDPT zero-shot (0.8821/0.8754). Across all sizes, Mitra consistently underperform (ACC <0.65<0.65), reflecting limited scalability.
Based on Dataset Width
Table 5: Performance variation by feature dimensionality (dataset width) across all benchmark suites. ACC = Accuracy; F1 = Weighted F1-score, averaged across datasets within each width category. Values are on a 0–1 scale (higher is better). Models are grouped by adaptation strategy. Formatting: 1st place ; 2nd place within each group.
Models Narrow (<10) Medium (10-100) Wide (>100)
Rank ACC F1 Rank ACC F1 Rank ACC F1
Baselines + Zero-Shot Inference
XGBoost 7.79 0.8199 0.8134 8.27 0.8482 0.8410 6.14 0.9101 0.9039
CatBoost 6.57 0.8124 0.8045 8.26 0.8441 0.8344 6.85 0.9127 0.9084
Random Forest 8.31 0.7981 0.7919 8.82 0.8410 0.8235 9.11 0.8987 0.8936
LightGBM 7.17 0.8107 0.8004 8.27 0.8458 0.8326 9.00 0.8957 0.8908
TabICL 6.01 0.8188 0.8098 5.62 0.8627 0.8549 7.93 0.8948 0.8936
OrionBiX 5.56 0.8089 0.8020 6.70 0.8510 0.8417 8.20 0.8859 0.8849
OrionMSP 3.16 0.8376 0.8295 4.98 0.8572 0.8478 7.00 0.8860 0.8837
TabPFN-2.0 6.08 0.8167 0.8071 4.93 0.8676 0.8589 7.44 0.9129 0.9111
TabPFN-2.5 4.63 0.8243 0.8120 3.92 0.8697 0.8609 4.20 0.9248 0.9240
Mitra 12.52 0.3769 0.2720 13.14 0.3886 0.2781 14.46 0.2521 0.1497
ContextTab 8.64 0.7965 0.7862 7.29 0.8441 0.8452 8.38 0.9008 0.8999
TabDPT 5.35 0.8243 0.8169 6.63 0.8566 0.8483 8.64 0.8845 0.8820
FineTune - Meta Learning
TabICL 3.66 0.7867 0.7779 3.48 0.8383 0.8338 3.86 0.8319 0.8667
OrionBiX 3.92 0.7956 0.7900 3.89 0.8316 0.8278 3.58 0.8429 0.8640
OrionMSP 1.90 0.8253 0.8190 2.21 0.8514 0.8457 2.86 0.8564 0.8737
TabPFN-2.0 2.73 0.8264 0.8171 2.37 0.8683 0.8579 2.28 0.9172 0.9533
Mitra 5.67 0.6646 0.6247 6.12 0.6004 0.5763 6.17 0.4090 0.3224
TabDPT 3.56 0.8129 0.8071 3.82 0.8405 0.8350 3.48 0.8400 0.8497
FineTune - Supervised FineTuning
TabICL 5.28 0.6609 0.6195 5.04 0.7507 0.7101 3.95 0.7757 0.7269
OrionBiX 4.03 0.6922 0.6576 4.33 0.7643 0.7294 3.75 0.7560 0.7046
OrionMSP 2.63 0.7540 0.7297 2.99 0.8025 0.7750 2.90 0.8148 0.7808
TabPFN-2.0 1.86 0.8258 0.8162 2.06 0.8578 0.8469 1.40 0.9346 0.9335
Mitra 5.81 0.5619 0.4478 6.09 0.5349 0.4268 5.60 0.3307 0.2436
TabDPT 2.50 0.8078 0.7948 2.88 0.8378 0.8294 3.40 0.8623 0.8587
FineTune – PEFT - Meta Learrning
TabICL 4.58 0.6375 0.6375 4.21 0.7601 0.7444 4.05 0.7714 0.7622
OrionBiX 2.51 0.7584 0.7584 2.92 0.8147 0.8057 2.50 0.8251 0.8243
OrionMSP 2.18 0.7836 0.7836 2.20 0.8254 0.8107 2.66 0.8293 0.8256
Mitra 4.15 0.6471 0.6471 4.83 0.6177 0.5650 4.61 0.4129 0.3338
TabDPT 2.22 0.7979 0.7979 2.32 0.8299 0.8209 2.94 0.8416 0.8376
FineTune - PEFT - Supervised FineTuning
TabICL 3.73 0.6196 0.5287 3.74 0.6596 0.5902 2.77 0.8760 0.8502
OrionBiX 3.81 0.6278 0.5444 4.08 0.6851 0.6177 4.00 0.6737 0.6166
OrionMSP 2.10 0.6795 0.6113 1.97 0.7477 0.6939 1.88 0.8042 0.7663
Mitra 4.28 0.5518 0.4481 4.78 0.5106 0.4108 4.77 0.2831 0.1874
TabDPT 1.72 0.7821 0.7631 2.02 0.8305 0.8201 1.55 0.8925 0.8901
Table 5 reveals that feature dimensionality strongly shapes model behaviour.
  • For narrow datasets (<<10 features), performance converges across models: OrionMSP zero-shot (0.8376/0.8295) slightly leads, followed by TabDPT and TabPFN, while classical baselines remain competitive.
  • On medium-width data (10–100 features), TabPFN dominates under both zero-shot (0.8376/0.8589) and meta-learning (0.8683/0.8579), outperforming all baselines and transformer variants.
  • The advantage of TFMs becomes most pronounced on wide feature spaces (>>100 features): TabPFN-2.0 achieves exceptional results under SFT (ACC=0.9346, F1=0.9335) and meta-learning (0.9172/0.9533), far exceeding other models. TabICL and OrionBiX degrade when fully fine-tuned, confirming that dense feature spaces accentuate overfitting. Traditional models (XGBoost) remain competitive but trail by 2–3 %.
Based on Class Imbalance
Table 6 compares balanced ($\geq 0.6$) and imbalanced (<0.6<0.6) scenarios.
  • Under balanced conditions, TabPFN (SFT) achieves the best rank (1.96) with ACC=0.8336, F1=0.8267, followed by OrionMSP (Meta Learning, rank 2.27).
  • When class ratios become highly skewed, meta-learning fine-tuning provides the greatest stability: OrionMSP achieves rank 2.16 (ACC=0.8735, F1=0.8636), with TabPFN close behind (rank=2.50, ACC=0.8784, F1=0.8664). These results confirm that episodic adaptation enhances calibration and minority-class reliability, while tree-based models lose 3–5% in accuracy under severe imbalance.
Table 6: Performance variation by class imbalance across all benchmark suites. ACC = Accuracy; F1 = Weighted F1-score, averaged within each imbalance category. Rank denotes the mean rank within each category (lower is better). Values are on a 0–1 scale (higher is better). Models are grouped by adaptation strategy. Formatting: 1st place ; 2nd place within each group.
Models / Strategy Balanced ($\geq 0.6$) Imbalanced (<0.6)
Rank ACC F1 Rank ACC F1
Baselines + Zero-Shot Inference
XGBoost 8.44 0.8175 0.8110 7.28 0.8853 0.8779
CatBoost 8.60 0.8076 0.8020 6.69 0.8780 0.8669
Random Forest 9.36 0.7983 0.7955 7.99 0.8736 0.8639
LightGBM 8.70 0.8071 0.7977 7.31 0.8769 0.8626
TabICL 5.81 0.8279 0.8233 6.00 0.8801 0.8692
OrionBiX 6.95 0.8096 0.8040 6.07 0.8782 0.8677
OrionMSP 5.20 0.8265 0.8202 3.96 0.8835 0.8725
TabPFN-2.0 4.73 0.8367 0.8309 6.23 0.8803 0.8691
TabPFN-2.5 3.81 0.8406 0.8348 4.49 0.8835 0.8727
Mitra 13.61 0.2763 0.1540 12.48 0.4824 0.3891
ContextTab 7.33 0.8152 0.8095 8.19 0.8722 0.8601
TabDPT 6.27 0.8233 0.8189 6.65 0.8793 0.8684
FineTune - Meta Learning
TabICL 3.78 0.7943 0.7884 3.58 0.8662 0.8547
OrionBiX 4.02 0.7947 0.7881 3.94 0.8622 0.8532
OrionMSP 2.27 0.8231 0.8183 2.16 0.8735 0.8636
TabPFN 2.31 0.8461 0.8407 2.50 0.8784 0.8664
Mitra 6.40 0.4699 0.4142 5.83 0.7885 0.7391
TabDPT 4.09 0.8066 0.8023 4.01 0.8650 0.8552
FineTune - Supervised FineTuning
TabICL 4.84 0.6708 0.6317 5.26 0.7794 0.7362
OrionBiX 4.11 0.6917 0.6550 4.32 0.7934 0.7590
OrionMSP 3.04 0.7467 0.7287 2.72 0.8307 0.7953
TabPFN 1.96 0.8336 0.8267 1.98 0.8709 0.8578
Mitra 6.40 0.3216 0.2037 5.57 0.7413 0.6415
TabDPT 3.02 0.8025 0.7949 2.57 0.8578 0.8464
FineTune – PEFT - Meta Learrning
TabICL 4.24 0.6530 0.6460 4.43 0.8026 0.7800
OrionBiX 2.73 0.7504 0.7436 2.84 0.8528 0.8420
OrionMSP 2.28 0.7714 0.7623 2.14 0.8608 0.8433
Mitra 5.02 0.4683 0.4173 4.21 0.8001 0.7524
TabDPT 2.30 0.7860 0.7801 2.33 0.8609 0.8511
FineTune - PEFT - Supervised FineTuning
TabICL 3.56 0.6003 0.5233 3.87 0.7223 0.6507
OrionBiX 4.02 0.5974 0.5164 3.96 0.7490 0.6875
OrionMSP 2.14 0.6753 0.6105 1.87 0.7915 0.7423
Mitra 4.97 0.3189 0.2171 4.21 0.7495 0.6497
TabDPT 1.76 0.7907 0.7798 2.10 0.8496 0.8346
Discussion : Scalability analysis underscores that TabPFN, OrionMSP, and TabDPT each occupy distinct strengths. TabPFN excels on small, medium, and high-dimensional data where Bayesian priors and strong supervision prevail; OrionMSP scales best with large datasets and imbalanced distributions due to meta-learning robustness; and TabDPT sustains near-top performance across all settings with minimal computational overhead. Classical baselines remain reliable but consistently trail the best TFMs by 2–4% in accuracy.
Key Takeaways:
  • DATASET SIZE: Small datasets favor the TabPFN family (TabPFN-2.5 zero-shot, TabPFN-2.0 meta-learning), with TabDPT as a close, efficient alternative; medium datasets favor TabPFN (meta & SFT); large datasets see XGBoost slightly leading in raw accuracy, with OrionMSP and TabDPT close behind.
  • Feature dimensionality: Wide datasets (>>100 features) favor TabPFN (SFT, ACC $0.92$, F1 $0.92$), confirming its strength for high-dimensional problems; medium-width data favor TabPFN (Meta Learning), while narrow data show minimal gaps between TFMs and baselines.
  • Class imbalance: Meta-learning fine-tuning yields the best robustness under imbalance, with OrionMSP (Rank-$2.16$, ACC = $0.8735$, F1 = $0.8636$) outperforming other TFMs; TabPFN (rank $2.50$) remains second.
  • Overall trend: TFMs scale smoothly with data size, width, and balance, maintaining >>2–4% advantage and >>3 rank improvement over traditional baselines.
6.2.3  Domain-Specific Analysis
To assess real-world applicability beyond general scalability patterns, we evaluate model performance across high-stakes domains including Medical and Finance (Table 7). Within domain-specific suites, different fine-tuning strategies yield distinct advantages depending on the domain.
Table 7: Domain-specific leaderboards for Medical and Finance datasets from the benchmark suites. Rank denotes the mean rank within each domain (lower is better). ACC = Accuracy; F1 = Weighted F1-score (0–1 scale, higher is better). Models are grouped by adaptation strategy. Formatting: 1st place ; 2nd place within each group.
Models / Strategy Medical Finance
Rank ACC F1 Rank ACC F1
Baselines + Zero-Shot Inference
XGBoost 7.32 0.7834 0.7669 7.62 0.7958 0.7885
RandomForest 7.30 0.7779 0.7752 8.46 0.8052 0.8001
CatBoost 7.48 0.7784 0.7594 6.75 0.8117 0.8015
LightGBM 6.36 0.7949 0.7614 7.32 0.8095 0.7974
TabICL 6.66 0.7819 0.7696 7.82 0.8125 0.7942
OrionBiX 4.98 0.7893 0.7759 6.46 0.8206 0.8125
OrionMSP 5.31 0.8045 0.7916 5.25 0.8158 0.8047
TabPFN-2.0 5.82 0.7984 0.7857 8.21 0.8094 0.7919
TabPFN-2.5 5.04 0.7965 0.7835 4.57 0.8169 0.8000
Mitra 12.00 0.3935 0.2863 15.14 0.5340 0.4250
ContextTab 6.40 0.8003 0.7881 10.85 0.8000 0.7830
TabDPT 7.96 0.7764 0.7641 9.46 0.8080 0.7960
FineTune - Supervised FineTuning
TabICL 4.64 0.7081 0.6565 6.50 0.6863 0.6335
OrionBiX 3.60 0.7400 0.6961 5.33 0.7389 0.7118
OrionMSP 2.24 0.7481 0.7116 3.00 0.7896 0.7523
TabPFN 1.86 0.8094 0.7958 2.00 0.8222 0.8058
Mitra 5.54 0.5600 0.4343 6.10 0.6823 0.5748
TabDPT 3.12 0.7717 0.7573 2.86 0.7922 0.7763
FineTune - Meta Learning
TabICL 3.40 0.7833 0.7695 3.43 0.8171 0.8019
OrionBiX 3.58 0.7623 0.7536 4.26 0.8106 0.8022
OrionMSP 2.20 0.7744 0.7645 2.26 0.8209 0.8089
TabPFN 2.28 0.8133 0.8000 3.00 0.8220 0.8056
Mitra 5.06 0.6499 0.5913 6.73 0.7559 0.7044
TabDPT 4.48 0.7535 0.7477 5.06 0.8021 0.7930
FineTune – PEFT - Meta Learrning
TabICL 4.19 0.7031 0.6873 5.07 0.7351 0.7118
OrionBiX 2.37 0.7628 0.7613 3.07 0.7958 0.7846
OrionMSP 2.02 0.7779 0.7658 2.57 0.7977 0.7842
Mitra 3.62 0.6539 0.6055 5.07 0.7492 0.7085
TabDPT 2.81 0.7390 0.7312 2.79 0.7934 0.7855
FineTune - PEFT - Supervised FineTuning
TabICL 3.68 0.5759 0.5044 3.57 0.7533 0.6833
OrionBiX 3.26 0.5926 0.5189 5.07 0.7279 0.6469
OrionMSP 2.02 0.6353 0.5767 1.93 0.7854 0.7272
Mitra 4.08 0.5557 0.4357 5.18 0.6665 0.5487
TabDPT 1.96 0.7680 0.7531 2.82 0.7782 0.7543
Medical Domain Performance : On medical datasets, TabPFN with supervised fine-tuning (SFT) achieves the best overall results (Rank = 1.86, ACC = 0.8094, F1 = 0.7958), followed by its meta-learning variant (Rank = 2.28, ACC = 0.8133, F1 = 0.8000). OrionMSP performs strongly in zero-shot (ACC = 0.8045, F1 = 0.7916) and meta-learning settings (Rank = 2.20, ACC = 0.7744, F1 = 0.7645), but exhibits minor degradation under full SFT. TabDPT (PEFT–SFT) also performs competitively (Rank = 1.96, ACC = 0.7680, F1 = 0.7531), achieving near-optimal results with reduced fine-tuning cost. In contrast, models such as TabICL and OrionBiX lose up to 7–10 % accuracy under SFT, confirming overfitting risk in small, noisy clinical datasets. Traditional baselines (XGBoost, LightGBM) remain competitive (ACC $\approx$ 0.79–0.80) but trail the best TFMs by 2–3 . Overall, the medical results demonstrate that meta-learning TFMs generalize more reliably under limited data and high noise.
Finance Domain Performance : Financial datasets are typically larger and more imbalanced, amplifying the value of episodic and parameter-efficient adaptation. OrionMSP under meta-learning attains the best overall rank (2.26) and strong accuracy (ACC = 0.8209, F1 = 0.8089), narrowly surpassing TabPFN (meta, Rank = 3.00, ACC = 0.8220, F1 = 0.8056) and its SFT counterpart (Rank = 2.00, ACC = 0.8222, F1 = 0.8058). TabDPT (PEFT–SFT) maintains competitive accuracy (Rank = 2.82, ACC = 0.7782, F1 = 0.7543), while OrionMSP (PEFT–SFT) achieves the strongest overall efficiency–accuracy balance (Rank = 1.93, ACC = 0.7854, F1 = 0.7272). Transformer models such as TabICL degrade sharply under SFT (ACC drop > 12 %), whereas classical baselines remain robust (ACC $\approx$ 0.81) but underperform in F1 due to poor minority calibration. The finance domain thus emphasizes meta-learning as the most stable strategy for large-scale, imbalanced tabular data.
Discussion : Across both domains, pretrained TFMs consistently outperform traditional baselines when appropriate fine-tuning strategies are used. TabPFN achieves the best reliability and calibration on small, noisy medical datasets, whereas OrionMSP dominates in large, high-imbalance financial contexts. TabDPT maintains strong performance in both settings through efficient PEFT adaptation, confirming that resource-efficient methods can match full SFT accuracy within 2 %. These trends highlight the need for strategy–domain alignment: Bayesian priors excel under data scarcity, while meta-learning transformers scale better with complex, heterogeneous features.
Key Takeaways:
  • Medical domain: TabPFN (SFT, rank $1.86$) achieves best overall accuracy, with its meta-learning variant (rank $2.28$) maintaining robustness; TabDPT (PEFT–SFT, rank $1.96$) offers strong low-cost alternatives.
  • Finance domain: OrionMSP (meta, rank $2.26$, ACC $\approx 0.82$) achieves top results, closely followed by TabPFN (meta/SFT, ranks 2.002.003.003.00); TabDPT (PEFT–SFT) remains efficient and competitive.
  • Cross-domain: TabPFN provides most consistent cross-domain performance; OrionMSP generalizes best to large-scale financial data; meta-learning and PEFT yield stable, domain-adaptive transfer.
6.3  Calibration Evaluation
Table 8: Probability calibration metrics across models and tuning strategies on the TALENT, OpenML-CC18, and TabZilla benchmark suites. ECE = Expected Calibration Error; MCE = Maximum Calibration Error; BS = Brier Score. All metrics range from 0–1 (lower is better). Values are averaged across datasets within each suite. Models are grouped by adaptation strategy. Formatting: 1st place ; 2nd place within each group.
Models / Strategy TALENT OpenML-CC18 TabZilla
ECE MCE BS ECE MCE BS ECE MCE BS
Zero-Shot Inference
TabICL 0.0219 0.2421 0.1533 0.0371 0.3404 0.1267 0.0369 0.2863 0.1301
OrionBiX 0.0324 0.2245 0.1787 0.0325 0.3230 0.1325 0.0385 0.2512 0.2419
OrionMSP 0.0219 0.2098 0.1589 0.0319 0.2902 0.1262 0.0310 0.2805 0.1243
TabPFN-2.0 0.0276 0.2470 0.1514 0.0375 0.2880 0.1253 0.0431 0.2745 0.1283
TabPFN-2.5 0.0266 0.2484 0.1483 0.0380 0.2945 0.1270 0.0530 0.2750 0.1272
Mitra 0.2238 0.3115 0.5291 0.2138 0.2952 0.5307 0.1946 0.2995 0.5733
ContextTab 0.02585 0.2525 0.1619 0.0352 0.3260 0.1353 0.0389 0.2815 0.1605
TabDPT 0.0308 0.2489 0.1586 0.0443 0.3160 0.1351 0.0435 0.2997 0.1227
FineTune - Meta Learning
TabICL 0.0355 0.2570 0.1671 0.0426 0.3387 0.1346 0.2033 0.4926 0.3732
OrionBiX 0.0409 0.2693 0.1841 0.0731 0.4164 0.1607 0.0744 0.3964 0.1416
OrionMSP 0.0340 0.2570 0.1656 0.0520 0.3458 0.1606 0.0365 0.2574 0.1250
TabPFN-2.0 0.0271 0.2440 0.1502 0.0346 0.2989 0.1178 0.0388 0.2636 0.1344
Mitra 0.0334 0.1912 0.3479 0.0533 0.2258 0.3932 0.0891 0.2086 0.4239
TabDPT 0.0562 0.3235 0.1791 0.0745 0.4178 0.1620 0.0664 0.2838 0.1591
FineTune - Supervised FineTuning
TabICL 0.0768 0.3219 0.2349 0.1156 0.3610 0.3039 0.1534 0.3280 0.4078
OrionBiX 0.0753 0.3204 0.2334 0.1131 0.3374 0.2923 0.1581 0.3801 0.3530
OrionMSP 0.0471 0.1886 0.2113 0.0550 0.3155 0.1995 0.0773 0.3321 0.2594
TabPFN-2.0 0.0287 0.2509 0.1580 0.0381 0.2957 0.1246 0.0469 0.2528 0.1454
Mitra 0.1723 0.2620 0.4677 0.1539 0.2269 0.4784 0.2023 0.2991 0.5254
TabDPT 0.0548 0.3019 0.1867 0.0497 0.3397 0.1443 0.0713 0.3183 0.1801
PEFT FineTune - Meta Learning
TabICL 0.0355 0.2570 0.1671 0.0426 0.3387 0.1346 0.2033 0.4926 0.3732
OrionBiX 0.0409 0.2693 0.1841 0.0731 0.4164 0.1607 0.0744 0.3964 0.1416
OrionMSP 0.0340 0.2570 0.1656 0.0520 0.3458 0.1606 0.0365 0.2574 0.1250
Mitra 0.0334 0.1912 0.3479 0.0533 0.2258 0.3932 0.0891 0.2086 0.4239
TabDPT 0.0562 0.3235 0.1791 0.0745 0.4178 0.1620 0.0664 0.2838 0.1591
PEFT FineTune - Supervised FineTuning
TabICL 0.0768 0.3219 0.2349 0.1156 0.3610 0.3039 0.1534 0.3280 0.4078
OrionBiX 0.0753 0.3204 0.2334 0.1131 0.3374 0.2923 0.1581 0.3801 0.3530
OrionMSP 0.0471 0.1886 0.2113 0.0550 0.3155 0.1995 0.0773 0.3321 0.2594
Mitra 0.1723 0.2620 0.4677 0.1539 0.2269 0.4784 0.2023 0.2991 0.5254
TabDPT 0.0548 0.3019 0.1867 0.0497 0.3397 0.1443 0.0713 0.3183 0.1801
Reliable probability calibration is essential for deploying tabular models in real-world decision-making systems. Calibration measures how well predicted probabilities align with observed outcomes, ensuring that model confidence reflects actual likelihoods. We evaluate Expected Calibration Error (ECE), Maximum Calibration Error (MCE), and Brier Score (BS) across all benchmark suites—TALENT, OpenML-CC18, and TabZilla—under different adaptation strategies. Lower values indicate better calibration. Table 8 summarizes the results and highlights how fine-tuning affects probabilistic reliability.
Zero-shot inference
[…]
Meta-learning fine-tuning
Episodic meta-learning largely preserves calibration quality. TabPFN achieves the best overall calibration across suites (ECE=0.0271–0.0388, BS=0.118–0.150), improving slightly over zero-shot results. OrionMSP remains competitive (ECE=0.0340–0.0520, BS=0.125–0.166), maintaining stable confidence estimates even after adaptation. In contrast, transformer-heavy models such as TabICL exhibit severe degradation on TabZilla (ECE rises from 0.0369 to 0.2033), showing that full episodic updates can destabilize latent priors when domain variability is high. Interestingly, Mitra improves calibration dramatically under meta-learning (ECE drops from $\approx 0.22$ to 0.030.030.050.05), although its Brier scores remain poor (>0.34>0.34), suggesting partial correction without overall reliability gains.
Supervised fine-tuning (SFT)
Full SFT substantially worsens calibration for most transformer models. TabICL’s ECE triples to quadruples (0.02–0.04 → 0.08–0.15) and its Brier Score deteriorates from $\approx 0.15$ to 0.230.230.410.41. OrionBiX follows a similar pattern. OrionMSP maintains moderate calibration (ECE=0.047–0.077, BS=0.21–0.26)—still degraded but markedly better than other transformers. TabPFN, however, preserves excellent reliability: across all suites, it records ECE=0.0287–0.0469, MCE<0.30, and BS=0.124–0.158, closely matching or outperforming its zero-shot performance. TabDPT also retains good calibration (ECE=0.0497–0.0713, BS=0.144–0.186). Mitra again performs worst (ECE > 0.15, BS > 0.47), underscoring structural limitations in its confidence estimation.
Discussion : Calibration analysis reveals a clear hierarchy of reliability. TabPFN consistently maintains among the lowest ECE and Brier scores across all adaptation strategies (ECE 0.027–0.047, BS 0.12–0.16), validating its Bayesian design and robustness to parameter updates. Meta-learning preserves calibration better than full SFT, confirming that episodic adaptation reinforces pretrained uncertainty priors. OrionMSP achieves strong calibration, especially under zero-shot and meta-learning, while transformer-centric models like TabICL and OrionBiX show pronounced drift once fully fine-tuned. Overall, TFMs maintain lower calibration error than ensemble baselines and degrade far less under fine-tuning, making them more trustworthy for risk-sensitive deployment.
Key Takeaways:
  • TabPFN maintains excellent calibration across all strategies (ECE 0.0270.0270.0470.047, BS 0.120.120.150.15), confirming its Bayesian reliability; TabDPT (SFT) and OrionMSP (zero-shot/meta-learning) rank next best (ECE 0.02190.02190.03650.0365).
  • Supervised fine-tuning degrades calibration for transformer models: TabICL’s ECE increases 3–4× and BS deteriorates sharply, while OrionMSP degrades moderately.
  • Zero-shot inference yields the best calibration overall, with OrionMSP and TabPFN showing ECE << 0.04 across all suites.
6.4  Fairness Evaluation
Table 9: Fairness metrics across models and tuning strategies on datasets with explicit demographic attributes. All metrics are absolute differences (0–1 range; lower is better, 0 = perfect equity). SPD = Statistical Parity Difference; EOD = Equalized Odds Difference; EOpD = Equalized Opportunity Difference. ACC = Accuracy; F1 = Weighted F1-score. Values are averaged across fairness evaluation datasets. Models are grouped by adaptation strategy. Sensitive features (e.g., race, gender, age) are manually specified per dataset. Formatting: 1st place ; 2nd place within each group.
Models / Strategy SPD EOD EOpD ACC F1
Zero-Shot Inference
TabICL 0.2900 0.3168 0.3278 0.8743 0.8680
OrionBiX 0.3328 0.2698 0.2855 0.8779 0.8727
OrionMSP 0.3380 0.2827 0.2983 0.8752 0.8706
TabPFN-2.0 0.3070 0.3114 0.3176 0.8708 0.8631
TabPFN-2.5 0.3067 0.3204 0.3144 0.8756 0.8679
Mitra 0.0193 0.0590 0.0982 0.6902 0.5847
ContextTab 0.2082 0.1860 0.1951 0.8580 0.8345
TabDPT 0.2990 0.3197 0.3338 0.8674 0.8582
FineTune - Meta Learning
TabICL 0.3140 0.3100 0.3332 0.8678 0.8603
OrionBiX 0.2021 0.1624 0.2010 0.8743 0.8692
OrionMSP 0.3038 0.2798 0.2907 0.8680 0.8650
TabPFN-2.0 0.3070 0.3115 0.3177 0.8733 0.8668
Mitra 0.2586 0.2748 0.3301 0.7967 0.7546
TabDPT 0.2648 0.3180 0.3080 0.8614 0.8503
FineTune - Supervised FineTuning
TabICL 0.0550 0.3225 0.3319 0.4584 0.3625
OrionBiX 0.1128 0.1761 0.1817 0.6180 0.5498
OrionMSP 0.1227 0.2208 0.2353 0.4753 0.4039
TabPFN-2.0 0.3070 0.3115 0.3177 0.8733 0.8668
Mitra 0.0161 0.0170 0.0555 0.7268 0.6277
TabDPT 0.3075 0.3153 0.3284 0.8529 0.8435
FineTune – PeFT - Meta Learning
TabICL 0.3140 0.3092 0.3332 0.8678 0.8602
OrionBiX 0.3149 0.2722 0.2875 0.8748 0.8717
OrionMSP 0.3038 0.2798 0.2907 0.867 0.8632
Mitra 0.2595 0.2391 0.3154 0.8104 0.7721
TabDPT 0.2778 0.2792 0.3051 0.8651 0.8556
FineTune – PeFT - SFT
TabICL 0.1572 0.4041 0.4089 0.4881 0.4269
OrionBiX 0.0349 0.1564 0.2234 0.6401 0.5759
OrionMSP 0.1132 0.2140 0.2420 0.5732 0.5080
Mitra 0.0029 0.0370 0.0740 0.7133 0.6028
TabDPT 0.2960 0.3190 0.3342 0.8595 0.8505
Fairness analysis evaluates whether model predictions exhibit systematic bias across demographic subgroups, complementing accuracy and calibration assessments by addressing equity in high-stakes applications such as lending, healthcare, and criminal justice. Unlike performance metrics that can be computed automatically, fairness requires explicit identification of sensitive attributes (e.g., race, gender). We use datasets with demographic labels—Adult Census Income, German Credit, and COMPAS Recidivism—to measure bias propagation and mitigation across strategies. Following standard fairness literature, we report Statistical Parity Difference (SPD), Equalized Odds Difference (EOD), and Equalized Opportunity Difference (EOpD). All metrics are absolute differences in the 0–1 range (lower = fairer). Table 9 summarizes fairness–accuracy trade-offs across zero-shot, meta-learning, and fine-tuning strategies.
Zero-shot inference.
A clear fairness–accuracy trade-off emerges. Mitra achieves the lowest bias (SPD =0.0193=0.0193, EOD =0.0590=0.0590, EOpD =0.0982=0.0982) but with very low predictive accuracy (ACC = 0.6902). OrionMSP and OrionBiX balance fairness and accuracy best, maintaining high accuracy (ACC = 0.875-0.878) and moderate fairness (EOD =0.27=0.270.280.28). TabPFN shows consistent behavior (ACC = 0.871, SPD = 0.307), while ContextTab ranks second in fairness (SPD =0.208=0.208) at slightly lower accuracy. These results suggest that transformer-based models preserve competitive fairness even without task-specific adaptation.
Meta-learning fine-tuning.
Episodic adaptation largely preserves fairness while sustaining accuracy. OrionBiX achieves the lowest overall bias (SPD = 0.2021, EOD = 0.1624, EOpD = 0.2010) while maintaining high accuracy (ACC = 0.8743). OrionMSP follows closely (EOD = 0.2798, EOpD = 0.2907),a close second with slightly higher bias but similar accuracy.
TabPFN achieves the highest accuracy (ACC = 0.8733) with slightly higher disparity (SPD $\approx$ 0.31). Mitra remains highly fair but again sacrifices accuracy. Overall, meta-learning fine-tuning maintains fairness levels similar to zero-shot inference while modestly improving accuracy consistency.
Supervised fine-tuning (SFT).
Full SFT produces mixed fairness outcomes. Mitra again records the lowest bias (SPD =0.0161=0.0161, EOD =0.0170=0.0170) but with only ACC = 0.7268. TabICL exhibits striking fairness gains (SPD drops from 0.29 to 0.055) yet loses over 40% in accuracy (ACC = 0.4584). OrionBiX improves fairness (EOD = 0.1761) but suffers large accuracy loss (ACC = 0.618). In contrast, TabPFN maintains both high accuracy (ACC = 0.8733) and moderate, stable fairness metrics across all strategies, confirming robustness. TabDPT ranks second in accuracy (ACC = 0.8529) with fairness levels comparable to TabPFN. These results underscore that SFT can reduce group disparity for some models but typically harms predictive reliability.
Parameter-efficient fine-tuning (PEFT).
PEFT variants show similar patterns while retaining higher stability. Under PEFT meta-learning, OrionBiX and OrionMSP achieve balanced fairness (EOD $\approx$ 0.27–0.28) and strong accuracy (ACC = 0.867–0.875), outperforming full SFT counterparts. Under PEFT SFT, Mitra again minimizes bias (SPD = 0.0029) but accuracy remains low (ACC = 0.7133). TabDPT delivers the best accuracy–fairness compromise (ACC = 0.8595, SPD $\approx$ 0.30). Overall, PEFT preserves fairness trends while mitigating the accuracy collapse observed in full fine-tuning.
Discussion.
Fairness evaluation confirms a fundamental accuracy–equity trade-off. Models achieving near-perfect demographic parity (Mitra, SPD <0.02<0.02) incur large accuracy losses, while TabPFN and OrionMSP consistently achieve balanced performance. Meta-learning proves most stable, preserving both fairness and accuracy across suites. SFT can occasionally improve parity for specific transformer architectures but typically at the expense of predictive fidelity. PEFT provides an effective compromise, maintaining high accuracy and moderate bias without significant fairness degradation.
Key Takeaways:
  • Accuracy–fairness trade-off: Mitra attains best fairness (SPD = 0.00290.00290.01930.0193) but suffers low accuracy (ACC = 0.69–0.73).
  • Balanced models: TabPFN and OrionMSP offer the best fairness–accuracy equilibrium (ACC = 0.87–0.88, EOD =0.27=0.270.290.29), outperforming other TFMs.
  • Effect of fine-tuning: SFT improves fairness for some models (e.g., TabICL SPD 0.055 vs 0.29 zero-shot) but drastically reduces accuracy; meta-learning preserves both aspects better.
  • PEFT stability: Parameter-efficient strategies maintain fairness trends while retaining high accuracy, offering a pragmatic balance for equitable deployment.
7  Discussion
Our comprehensive analysis across three benchmark suites reveals clear patterns that guide model selection, adaptation strategy, and deployment trade-offs.
7.1  Fine-Tuning Strategy Recommendations
Fine-tuning strategy choice is critical and highly model-dependent.
  • Zero-shot inference achieves excellent calibration (ECE <<0.04 for top models) and competitive predictive performance. It is ideal when computational resources are limited or calibration reliability is a priority. TabPFN-2.5 attains the highest overall zero-shot standing (Rank = 4.13), establishing a new state-of-the-art for inference-only tabular tasks. OrionMSP follows closely, achieving strong accuracy on large datasets (ACC = 0.8722–0.8821) and ranking among the top configurations on OpenML-CC18 and TabZilla without requiring additional training.
  • Meta-learning fine-tuning provides the best balance between adaptation and generalization. It improves rank stability for models such as OrionMSP (aggregate rank $2.26$ in Table 3) while preserving calibration (ECE <<0.06). Meta-learning also excels in high-imbalance settings (OrionMSP rank $2.16$, TabPFN rank $2.50$ in Table 6), confirming its cross-domain robustness. However, instability can occur for some models such as TabICL on TabZilla (ECE increases to $0.2033$), requiring careful validation.
  • Supervised fine-tuning (SFT) delivers the strongest results for TabPFN (ranks 1.831.831.891.89, ACC up to $0.9346$ on wide datasets). In contrast, SFT causes substantial degradation for TabICL and OrionBiX: for instance, TabICL’s accuracy on TabZilla falls from $0.8734$ to $0.5670$, indicating severe overfitting. Calibration also deteriorates (ECE 0.080.080.150.15 vs. 0.020.020.040.04 zero-shot). SFT should therefore be used selectively—beneficial for TabPFN, acceptable for TabDPT and OrionMSP, but risky for TabICL and OrionBiX.
  • Parameter-efficient fine-tuning (PEFT) attains nearly full fine-tuning accuracy while greatly reducing computational cost. TabDPT achieves high performance ($\text{ACC}\approx 0.85$, $\text{F1}\approx 0.84$) across all benchmarks, and OrionMSP maintains competitive accuracy ($\text{ACC}\approx 0.82$$0.86$, $\text{F1}\approx 0.80$$0.85$) with strong cross-dataset generalization.
7.2  Mechanisms Behind Fine-Tuning Performance
The varying outcomes of different fine-tuning strategies can be explained by the underlying adaptation mechanisms.
Why SFT Underperforms for TabICL and OrionBiX?
Full supervised fine-tuning updates all model parameters, erasing pretrained representations—especially detrimental on smaller datasets. The performance collapse of TabICL on TabZilla (ACC $0.8734\rightarrow 0.5670$) exemplifies overfitting and representation drift, whereas optimization distorts pretrained priors. Both TabICL and OrionBiX rely on high-capacity transformer layers optimized for diverse pretraining data; constraining them to a narrow distribution leads to loss of generality and inflated variance in predictions.
Why Meta-Learning Preserves Generalization?
Episodic training in OrionMSP and TabPFN aligns closely with the few-shot adaptation process. Each meta-episode simulates the support–query dynamics of deployment, encouraging transfer of cross-task features rather than memorization. This prevents catastrophic forgetting and reinforces pretrained inductive biases. The improvement in rank stability for OrionMSP (from 2.902.903.843.84 in zero-shot to 1.731.732.822.82 in meta-learning) confirms that episodic optimization promotes stable generalization.
When Fine-Tuning is Beneficial?
Fine-tuning is most useful when:
  1. 1.
    Datasets are sufficiently large (>>10K samples), mitigating overfitting.
  2. 2.
    Domain distribution shifts significantly from pretraining data (e.g., medical, finance).
  3. 3.
    Architecture maintains calibration under parameter updates (as in TabPFN’s Bayesian prior design).
For small or balanced datasets, zero-shot or meta-learning is more reliable and cost-efficient. Parameter-efficient fine-tuning (PEFT) provides a strong middle ground, retaining $\approx$95% of full SFT accuracy while reducing memory.
7.3  Calibration Insights
[…]
7.4  Model Selection Guidelines
Model selection should be guided by dataset scale, balance, feature dimensionality, and available computational resources. All recommendations below are derived from the common subset of datasets evaluated across all models to ensure consistent statistical comparison. Reported values combine mean accuracy (ACC), weighted F1-score (F1), and mean rank (based on accuracy) as presented in Tables 36.
General Guidelines:
When dataset characteristics are unknown, zero-shot inference with OrionMSP or meta-learning fine-tuning with TabPFN serve as the most reliable defaults. Both models consistently achieve top-three average ranks across all benchmark suites while maintaining strong calibration and fairness. In general:
  • TabPFN provides the most consistent accuracy and F1 across diverse data regimes, excelling in calibration and few-shot generalization.
  • OrionMSP scales best on large, complex, and imbalanced datasets, particularly under meta-learning.
Specific Regime Recommendations.
  • On small datasets (<<1K samples), TabPFN-2.5 achieves the highest zero-shot performance (ACC = 0.8349, F1 = 0.8194), slightly ahead of TabDPT (0.8333/0.8271) and boosted-tree baselines. Under meta-learning, TabPFN-2.0 also leads (0.8336/0.8256), while OrionMSP (PEFT–Meta) provides strong performance at lower fine-tuning cost (0.8275/0.8213). Fine-tuning TabICL or OrionBiX on small datasets still causes sharp degradation (ACC drops $15\%$), revealing overfitting risk.
  • Medium datasets (1K–10K samples): TabPFN with meta-learning achieves the highest mean accuracy (ACC = 0.8638, F1 = 0.8548, rank $\approx 2.0$), followed closely by its supervised fine-tuning variant (ACC = 0.8580, F1 = 0.8485). When computational constraints are present, TabDPT (PEFT–Meta) maintains competitive accuracy (ACC = 0.8121–0.8332) with minimal resource overhead. Zero-shot inference remains a practical option for quick deployment, trailing fine-tuned variants by only about 2233% in accuracy.
  • Large datasets (>>10K samples): OrionMSP in zero-shot mode delivers the highest mean performance among TFMs (ACC = 0.8843, F1 = 0.8768, rank $\approx 2.0$). TabDPT (zero-shot or PEFT–SFT) achieves comparable accuracy (ACC = 0.8496–0.8831) while offering lower adaptation cost. Classical models such as XGBoost and CatBoost reach strong absolute accuracies (ACC = 0.8969, F1 = 0.8920) but exhibit poorer calibration and higher cross-dataset variance. TFMs are therefore more advantageous when generalization consistency or uncertainty estimation is prioritized.
  • Balanced datasets: TabPFN under supervised fine-tuning yields the best average rank (1.96) with ACC = 0.8336, F1 = 0.8267, followed by OrionMSP under meta-learning (rank $2.27$). Balanced data favor Bayesian or episodically fine-tuned architectures, which maximize calibration and stability.
  • Imbalanced datasets: Meta-learning strategies provide the most robust adaptation under label imbalance. OrionMSP achieves the best overall performance (rank = 2.16, ACC = 0.8735, F1 = 0.8636), with TabPFN ranking second (rank = 2.50, ACC = 0.8784, F1 = 0.8664). These results confirm that episodic training improves minority-class calibration. In contrast, tree-based baselines lose roughly 3355% in F1 when class ratios fall below $0.6$.
  • Low-to-medium feature dimensionality (<<100 features): Model performance converges across methods (ACC = 0.84–0.86), with TabPFN (Meta Learning) and OrionMSP (Zero-Shot) performing most consistently. In this regime, TFMs primarily outperform classical methods in calibration and rank stability rather than raw accuracy.
  • High-dimensional datasets (>>100 features): TabPFN under supervised fine-tuning achieves the highest overall scores (ACC = 0.9346, F1 = 0.9335), significantly surpassing other models. TabDPT and OrionMSP maintain strong accuracy (ACC $\approx$ 0.84–0.89) while offering superior efficiency. Architectures such as TabICL and OrionBiX exhibit sharp performance drops under full fine-tuning, confirming their susceptibility to overfitting in wide-feature settings.
Key Takeaways
  • For inference-only deployment, use OrionMSP (Zero-Shot) or TabPFN (Zero-Shot) for balanced accuracy and calibration.
  • For adaptive fine-tuning, prefer TabPFN (SFT/Meta) on small–to–medium datasets and OrionMSP (Meta) on large-scale or imbalanced data.
7.5  Domain-Specific Guidance
Medical domain: TabPFN achieves the strongest performance in this domain, with its supervised fine-tuning configuration attaining Rank = 1.86 and ACC = 0.8094, closely followed by the meta-learning variant (Rank = 2.28, ACC = 0.8133). TabDPT under PEFT–SFT (Rank = 1.96) provides competitive accuracy at substantially lower computational cost. Transformer-based architectures such as TabICL and OrionBiX exhibit notable degradation under full fine-tuning, reflecting sensitivity to small and noisy clinical datasets. Overall, Bayesian and episodically trained models generalize more reliably in data-scarce medical settings.
Finance domain: OrionMSP under meta-learning achieves strong results (Rank = 2.26, ACC = 0.8209, F1 = 0.8089), while TabPFN performs exceptionally well in supervised and meta-learning modes (Ranks = 2.00--3.00, ACC = 0.8220--0.8222). TabDPT remains competitive (PEFT--SFT, Rank = 2.82) and offers an efficient alternative with only a minor loss in accuracy. These outcomes confirm that episodic and parameter-efficient adaptation strategies are particularly effective for large-scale, class-imbalanced financial data, where they improve both calibration and minority-class performance.
7.6  Practical Value of Tabular Foundation Models
Tabular Foundation Models (TFMs) deliver consistent but moderate absolute accuracy improvements—typically 2244 percentage points—over strong classic baselines such as XGBoost, CatBoost, and LightGBM. Across comparable settings, TFMs achieve mean accuracies around 0.850.850.880.88 versus 0.830.830.860.86 for boosted trees. However, their primary advantage lies in reliability and robustness rather than raw accuracy.
Key strengths include: (I) ZERO-SHOT ADAPTABILITY: The TabPFN family and OrionMSP provide top-tier zero-shot performance. TabPFN-2.5 attains the best aggregate zero-shot rank (4.13) with ACC $\approx\textsc{0.85}$0.88, while OrionMSP remains highly competitive (rank 4.61, ACC = 0.84550.8821). Both outperform boosted-tree baselines by roughly 1.5–2.5 percentage points in accuracy without any task-specific training. (II) Calibration: TabPFN maintains low calibration error (ECE <0.05<0.05) across both zero-shot and fine-tuned regimes, outperforming ensemble baselines whose ECE often exceeds $0.08$. (III) Fairness stability: TFMs achieve moderate statistical parity differences (SPD $\approx 0.28$$0.30$) while sustaining high accuracy (>0.85>0.85), indicating balanced performance across sensitive groups.
Consequently, TFMs should be regarded as reliable, general-purpose learners well suited for rapid deployment, few-shot adaptation, and risk-sensitive applications where calibration and fairness are critical. For large, well-labeled, single-domain problems, gradient-boosted trees may still be preferable for simplicity and computational efficiency, but TFMs offer distinct advantages in uncertainty estimation, cross-domain generalization, and calibration consistency.
7.7  Limitations
While TabTune provides broad support for evaluating and fine-tuning TFMs, several limitations remain.
Task Scope: Current support is limited to binary and multi-class classification tasks; regression and multi-label extensions are planned for future releases.
Model Constraints:
  • Dataset size: TabPFN performs optimally on datasets with fewer than 10,00010{,}000 samples; larger datasets may significantly increase inference latency or exceed memory limits.
  • Dimensionality: All models can degrade when the feature space exceeds $\sim 1000$ dimensions without preprocessing or feature selection.
  • PEFT compatibility: Parameter-efficient fine-tuning is fully supported for OrionMSP and TabDPT, but only partially implemented for TabPFN and ContextTab.
  • ContextTab evaluation: Due to excessive GPU memory requirements, ContextTab was evaluated only on smaller benchmark subsets.
Evaluation and Analysis:
  • Manual fairness specification: Sensitive attributes are manually defined per dataset, introducing potential bias in fairness evaluation.
  • Calibration interpretation: Optimal threshold tuning remains domain-dependent and may vary across metrics or class distributions.
  • Beyond-accuracy metrics: Advanced interpretability, reliability, and uncertainty quantification metrics are not yet fully integrated into the current release.
8  Conclusion
We introduce TabTune, a unified Python library that standardizes the workflow for tabular foundation models (TFMs). By abstracting model-specific preprocessing, consolidating adaptation strategies, and integrating evaluation modules for performance, calibration, and fairness, TabTune addresses the fragmentation that has limited the broader adoption of TFMs. Inspired by AutoML, TabTune focuses on managing and optimizing pretrained TFMs rather than building models from scratch, offering a coherent framework for efficient adaptation and assessment.
We conducted a comprehensive study on the performance of tabular foundation models (TFMs) and the impact of fine-tuning strategies on their effectiveness. The empirical analysis demonstrates that the choice of adaptation strategy is highly model-dependent. TabPFN performs optimally under supervised fine-tuning or meta-learning, OrionMSP exhibits strong generalization and robustness particularly with meta-learning, and TabDPT using parameter-efficient fine-tuning (PEFT) achieves a favorable balance between predictive accuracy and computational efficiency. Beyond raw performance, our findings emphasize the reliability and fairness of TFMs. TabPFN consistently attains excellent calibration, while both OrionMSP and TabPFN maintain a strong fairness–accuracy balance, reinforcing their potential for responsible and trustworthy deployment in high-stakes domains.
9  Future Work
As an extensible and reproducible platform for TFM research, TabTune continues to evolve to support emerging methodologies and broader evaluation dimensions. Future work will focus on integrating more advanced fine-tuning and meta-learning approaches to enhance few-shot and domain adaptation capabilities. In parallel, we aim to strengthen the evaluation framework by incorporating modules for advanced fairness assessment, interpretability, and uncertainty quantification. Additionally, we plan to extend task coverage beyond binary and multi-class classification to include regression, multi-label learning, and time-series modeling. By standardizing the TFM workflow, TabTune enables practitioners to transition efficiently from experimentation to reliable deployment, fostering evidence-based model selection and accelerating the responsible adoption of foundation models in structured data domains.
10  Appendix A: Default Hyperparameters
This appendix provides detailed default hyperparameters for all tuning strategies supported in TabTune. These values are optimized for general-purpose use and can be overridden via the tuning_params configuration dictionary.
10.1  Zero-Shot Inference
Zero-shot inference uses pretrained model weights without any parameter updates. No training hyperparameters are required for this strategy. The model receives training samples as context during inference and performs predictions through forward passes alone. However, each model has specific inference parameters that control prediction behavior:
Parameter Descriptions:
  • n_estimators / n_ensembles: Number of ensemble runs with different configurations for robust predictions
  • softmax_temperature / temperature: Scaling factor for logits before softmax (lower values yield sharper predictions)
  • context_size: Number of training samples used as context for each test prediction
  • average_before_softmax: Whether to average logits or probabilities across ensemble runs
  • safety_factor: Memory safety factor for batch size estimation (TabICL/OrionMSP/OrionBiX)
  • offload: Whether to offload intermediate computations to CPU for memory efficiency
  • use_amp: Automatic mixed precision for faster inference with reduced memory
Table 10: Default inference hyperparameters for zero-shot inference across models. Values shown are default settings used when no custom parameters are specified. All parameters can be customized via model-specific configuration dictionaries.
Model Inference Parameters
TabPFN n_estimators=8, softmax_temperature=0.9, average_before_softmax=False
TabICL Inference config: min_batch_size=1, safety_factor=0.8, offload=auto (COL), False (ROW/ICL), use_amp=True
OrionMSP Inference config: min_batch_size=1, safety_factor=0.8, offload=auto (COL), False (ROW/ICL), use_amp=True
OrionBiX Inference config: min_batch_size=1, safety_factor=0.8, offload=auto (COL), False (ROW/ICL), use_amp=True
TabDPT n_ensembles=8, temperature=0.8, context_size=512, permute_classes=True
Mitra d_model=64, num_heads=4, num_layers=2, use_synthetic_prior=True
ContextTab (No specific inference parameters)
10.2  Supervised Fine-Tuning (SFT)
Supervised Fine-Tuning (SFT) is a standard adaptation approach where the entire pretrained model is updated using gradient descent on task-specific labeled data. Unlike meta-learning, SFT treats the target dataset as a single supervised learning task, training the model end-to-end to minimize cross-entropy loss over multiple epochs.
Key Characteristics:
  • Training Paradigm: Standard supervised learning with full-batch or mini-batch optimization
  • Parameter Updates: All model parameters are updated during training
  • Objective: Minimize classification loss on the target task’s training set
  • Use Cases: Optimal for large datasets (>10K samples) where task-specific optimization is beneficial
  • Advantages: Can achieve highest task-specific accuracy, full model capacity utilization
  • Considerations: Higher memory requirements, risk of overfitting on small datasets, may reduce generalization to new tasks
General Hyperparameter Guidelines:
  • Learning Rate: Typically ranges from 1e-5 to 1e-4, with lower rates preserving pretrained knowledge
  • Epochs: 3-25 epochs depending on dataset size and convergence behavior
  • Batch Size: 16-128, adjusted based on dataset size and memory constraints
  • Regularization: Weight decay (1e-4 to 1e-5) and learning rate warmup help stabilize training
  • Optimizer: Adam or AdamW are standard choices for tabular foundation models
Table 11: Default hyperparameters for Supervised Fine-Tuning (SFT) across models. SFT updates all model parameters using gradient descent on task-specific labeled data. All parameters can be overridden via the tuning_params configuration dictionary. Optimal for large datasets (>>10K samples).
Model Key Parameters
TabPFN epochs=25, learning_rate=1e-5, max_episode_size=len(X), query_set_ratio=0.3, weight_decay=1e-4, optimizer=AdamW
TabICL epochs=5, learning_rate=1e-5, batch_size=16, optimizer=Adam
OrionMSP epochs=5, learning_rate=1e-5, batch_size=16, optimizer=Adam
OrionBiX epochs=5, learning_rate=1e-5, batch_size=16, optimizer=Adam
Mitra epochs=5, learning_rate=1e-5, batch_size=128, weight_decay=1e-4, warmup_epochs=1, optimizer=Adam
TabDPT epochs=5, learning_rate=2e-5, batch_size=32, weight_decay=1e-4, warmup_epochs=1, optimizer=Adam
ContextTab epochs=5, learning_rate=1e-4, batch_size=128, optimizer=Adam
10.3  Meta-Learning Fine-Tuning
Meta-learning fine-tuning, also known as episodic fine-tuning or few-shot adaptation, mimics the few-shot learning scenario during training. The model is trained using episodic data sampling, where each episode consists of a support set (context) and a query set (target predictions). This approach trains the model to quickly adapt to new tasks from limited examples, improving generalization.
Key Characteristics:
  • Training Paradigm: Episodic learning with support-query splits
  • Episode Structure: Each training step samples support and query sets from the target dataset
  • Objective: Learn to make accurate predictions on query sets given support set context
  • Use Cases: Ideal for small to medium datasets (<10K samples), maintains generalization to new tasks
  • Advantages: Better generalization, mimics actual inference scenarios for in-context learning models
  • Considerations: Requires careful episode sampling, support/query size tuning, more epochs needed
General Hyperparameter Guidelines:
  • Learning Rate: Typically lower than SFT (2e-6 to 1e-5) to preserve pretrained representations
  • Support Size: 32-512 samples per episode, depending on dataset size and model architecture
  • Query Size: Usually 20-50% of support size, determines episode difficulty
  • Episodes per Epoch: 100-1000 episodes ensure diverse task exposure
  • Epochs: 3-5 epochs often sufficient due to episodic diversity
  • Batch Size: 1-8 episodes per batch for memory-efficient training
Episode Filtering: Episodes where query classes are absent from the support set are automatically filtered to maintain class consistency. This ensures the model always has context examples for all classes it must predict.
Table 12: Default hyperparameters for Meta-Learning Fine-Tuning across models. Meta-learning uses episodic training with support-query splits. Episodes where query classes are absent from support set are automatically filtered. All parameters can be customized via tuning_params. Ideal for small to medium datasets (<<10K samples).
Model Key Parameters
TabPFN epochs=3, learning_rate=1e-5, batch_size=256, optimizer=AdamW
TabICL epochs=5, learning_rate=2e-6, support_size=48, query_size=32, n_episodes=1000, optimizer=Adam
OrionMSP epochs=5, learning_rate=2e-6, support_size=48, query_size=32, n_episodes=1000, optimizer=Adam
OrionBiX epochs=5, learning_rate=2e-6, support_size=48, query_size=32, n_episodes=1000, optimizer=Adam
Mitra epochs=3, learning_rate=1e-5, batch_size=4, support_size=128, query_size=128, steps_per_epoch=50, optimizer=Adam
TabDPT epochs=5, learning_rate=1e-5, batch_size=8, support_size=512, query_size=256, steps_per_epoch=100, optimizer=Adam
10.4  Parameter-Efficient Fine-Tuning (PEFT)
Parameter-Efficient Fine-Tuning (PEFT) reduces memory and computational requirements by updating only a small subset of model parameters through low-rank adaptation techniques. PEFT can be combined with either SFT or meta-learning strategies, providing efficient alternatives to full parameter updates.
Key Characteristics:
  • Training Paradigm: Updates only adapter parameters, base model weights remain frozen
  • Memory Efficiency: Reduces fine-tuning memory by 60-80% compared to full fine-tuning
  • Parameter Updates: Only low-rank adapter matrices are trained, not the full model
  • Use Cases: Memory-constrained environments, rapid experimentation, multiple task adaptations
  • Advantages: Fast training, lower memory footprint, easy task switching (swap adapters)
  • Considerations: May achieve slightly lower accuracy than full fine-tuning, model compatibility varies
Low-Rank Adaptation (LoRA) Mechanism:
LoRA introduces trainable low-rank matrices that approximate full weight updates. For a linear layer with weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA adds trainable matrices $A\in\mathbb{R}^{m\times r}$ and $B\in\mathbb{R}^{r\times n}$ where $r\ll\min(m,n)$ is the rank. The forward pass becomes:
$h=Wx+\frac{\alpha}{r}(BA)x$
where $\alpha$ is a scaling factor and $BA$ is the low-rank update. Only $A$ and $B$ are updated during training, while $W$ remains frozen.
General Hyperparameter Guidelines:
  • Rank ($r$): Controls adapter capacity; typically 4-16, with 8 providing good balance
  • LoRA Alpha ($\alpha$): Scaling factor; typically $\alpha=2r$ (e.g., $\alpha=16$ for $r=8$)
  • LoRA Dropout: 0.05-0.2 for regularization, prevents adapter overfitting
  • Target Modules: Attention projections, feed-forward layers, and encoders are common targets
  • Base Strategy: Uses same epochs, learning rate, and batch size as SFT or meta-learning
PEFT-SFT vs PEFT Meta-Learning: PEFT can be applied to either strategy:
  • PEFT-SFT: Combines full-batch supervised learning with LoRA adapters; best for large datasets
  • PEFT Meta-Learning: Combines episodic training with LoRA adapters; maintains generalization benefits
PEFT-SFT: Uses SFT hyperparameters (Table 11) with LoRA adapters applied. Compatible with: [TabICL, OrionMSP, OrionBiX, TabDPT, Mitra].
PEFT Meta-Learning: Uses Meta-Learning hyperparameters (Table 12) with LoRA adapters applied. Compatible with: [TabICL, OrionMSP, OrionBiX, TabDPT, Mitra].
Note: TabPFN and ContextTab have limited or experimental PEFT support. TabPFN PEFT is unstable due to architectural constraints with batched inference, and ContextTab PEFT may conflict with its embedding pipeline. These models automatically fall back to base fine-tuning when PEFT is requested.
Table 13: Default LoRA (Low-Rank Adaptation) configuration for Parameter-Efficient Fine-Tuning (PEFT). PEFT reduces memory by 60-80% compared to full fine-tuning. These values apply to both PEFT-SFT and PEFT Meta-Learning strategies. All parameters can be customized via peft_config dictionary.
Parameter Value
rank ($r$) 8
LoRA alpha ($\alpha$) 16
LoRA dropout 0.05
Target modules Model-specific (attention projections, encoders, transformers)
10.5  Additional Notes
All training strategies use CrossEntropyLoss for classification tasks. For meta-learning, episodes with query classes absent from support sets are automatically filtered to maintain consistency. Sampling strategies follow the guidelines in Section 4.2.3. Gradient clipping and learning rate schedulers (e.g., warmup) are applied where specified in the configuration above.
11  Appendix B: Experimental Setup
11.1  Hardware and Software Configuration
All experiments were executed on NVIDIA L40S GPUs. For some experiments we get OOM issues for those cases we use NVIDIA H200 GPUs.
11.2  Data Splitting and Reproducibility
All benchmarks follow their respective standardized dataset splits. Random seeds are fixed for reproducibility across experiments. Cross-validation is performed where specified by the benchmark protocols (TALENT uses auto-discovered splits, OpenML-CC18 uses predefined splits, TabZilla follows its standard evaluation protocol).
11.3  Dataset Statistics
The following subsections provide comprehensive statistics for all datasets used in our evaluation across different benchmark suites. Each table includes dataset names, domains, sample counts, feature counts, number of classes, and task types (binary classification or multiclass classification).
OpenML-CC18 Benchmark Datasets
Table LABEL:tab:datasets_openml lists all datasets from the OpenML-CC18 benchmark suite used in our evaluation.
TALENT Benchmark Datasets
Table LABEL:tab:datasets_talent lists all datasets from the TALENT benchmark suite used in our evaluation.
TabZilla Benchmark Datasets
Table LABEL:tab:datasets_tabzilla lists all datasets from the TabZilla benchmark suite. TabZilla uses OpenML dataset IDs, and these datasets are specifically selected for evaluating neural network performance on tabular data.
Fairness Evaluation Datasets
Fairness evaluation requires datasets with sensitive features (e.g., demographic attributes). Table LABEL:tab:datasets_fairness lists commonly used datasets for fairness evaluation. These datasets contain demographic or other sensitive attributes that enable fairness metric computation (Statistical Parity Difference, Equalized Odds Difference, Equalized Opportunity Difference).
11.4  Evaluation Metrics
All experiments report the following metrics:
  • Performance: Accuracy, AUC-ROC, weighted F1-score
  • Calibration: Expected Calibration Error (ECE), Maximum Calibration Error (MCE), Brier score
  • Fairness: Statistical Parity Difference (SPD), Equalized Odds Difference (EOD), Equalized Opportunity Difference (EOpD)
All metrics are computed using standard implementations and follow established benchmark protocols for each suite.
Table 14: OpenML-CC18 benchmark datasets (72 datasets).
Benchmark Dataset Name Domain Samples Features Classes Task Type Used In Experimentation
OpenML OpenML-ID-3 Other 3196 37 2 binclass Yes
OpenML OpenML-ID-6 Handwriting 20000 17 26 multiclass No
OpenML OpenML-ID-11 Other 625 5 3 multiclass Yes
OpenML OpenML-ID-12 Other 2000 217 10 multiclass Yes
OpenML OpenML-ID-14 Other 2000 77 10 multiclass Yes
OpenML OpenML-ID-15 Healthcare 699 10 2 binclass Yes
OpenML OpenML-ID-16 Other 2000 65 10 multiclass Yes
OpenML OpenML-ID-18 Other 2000 7 10 multiclass Yes
OpenML OpenML-ID-22 Other 2000 48 10 multiclass Yes
OpenML OpenML-ID-23 Healthcare 1473 10 3 multiclass Yes
OpenML OpenML-ID-28 Handwriting 5620 65 10 multiclass Yes
OpenML OpenML-ID-29 Finance 690 16 2 binclass Yes
OpenML OpenML-ID-31 Finance 1000 21 2 binclass Yes
OpenML OpenML-ID-32 Handwriting 10992 17 10 multiclass Yes
OpenML OpenML-ID-37 Healthcare 768 9 2 binclass Yes
OpenML OpenML-ID-38 Healthcare 3772 30 2 binclass Yes
OpenML OpenML-ID-44 Other 4601 58 2 binclass Yes
OpenML OpenML-ID-46 Other 3190 61 3 multiclass Yes
OpenML OpenML-ID-50 Other 958 10 2 binclass Yes
OpenML OpenML-ID-54 Other 846 19 4 multiclass Yes
OpenML OpenML-ID-151 Other 45312 9 2 binclass Yes
OpenML OpenML-ID-182 Other 6430 37 6 multiclass Yes
OpenML OpenML-ID-188 Other 736 20 5 multiclass Yes
OpenML OpenML-ID-300 Other 7797 618 26 multiclass No
OpenML OpenML-ID-307 Other 990 13 11 multiclass No
OpenML OpenML-ID-458 Other 841 71 4 multiclass Yes
OpenML OpenML-ID-469 Healthcare 797 5 6 multiclass Yes
OpenML OpenML-ID-554 Other 70000 785 10 multiclass Yes
OpenML OpenML-ID-1049 Other 1458 38 2 binclass Yes
OpenML OpenML-ID-1050 Other 1563 38 2 binclass Yes
OpenML OpenML-ID-1053 Other 10885 22 2 binclass Yes
OpenML OpenML-ID-1063 Other 522 22 2 binclass Yes
OpenML OpenML-ID-1067 Other 2109 22 2 binclass Yes
OpenML OpenML-ID-1068 Other 1109 22 2 binclass Yes
OpenML OpenML-ID-1461 Finance 45211 17 2 binclass Yes
OpenML OpenML-ID-1462 Finance 1372 5 2 binclass Yes
OpenML OpenML-ID-1464 Healthcare 748 5 2 binclass Yes
OpenML OpenML-ID-1468 Other 1080 857 9 multiclass Yes
OpenML OpenML-ID-1475 Other 6118 52 6 multiclass Yes
OpenML OpenML-ID-1478 Other 10299 562 6 multiclass Yes
OpenML OpenML-ID-1480 Healthcare 583 11 2 binclass Yes
OpenML OpenML-ID-1485 Other 2600 501 2 binclass Yes
OpenML OpenML-ID-1486 Other 34465 119 2 binclass No
OpenML OpenML-ID-1487 Other 2534 73 2 binclass Yes
OpenML OpenML-ID-1489 Other 5404 6 2 binclass Yes
OpenML OpenML-ID-1494 Other 1055 42 2 binclass Yes
OpenML OpenML-ID-1497 Other 5456 25 4 multiclass Yes
OpenML OpenML-ID-1501 Other 1593 257 10 multiclass Yes
OpenML OpenML-ID-1510 Other 569 31 2 binclass Yes
OpenML OpenML-ID-1590 Other 48842 15 2 binclass Yes
OpenML OpenML-ID-4134 Other 3751 1777 2 binclass No
OpenML OpenML-ID-4534 Other 11055 31 2 binclass Yes
OpenML OpenML-ID-4538 Other 9873 33 5 multiclass Yes
OpenML OpenML-ID-6332 Other 540 40 2 binclass Yes
OpenML OpenML-ID-23381 Retail 500 13 2 binclass Yes
OpenML OpenML-ID-23517 Other 96320 22 2 binclass No
OpenML OpenML-ID-40499 Other 5500 41 11 multiclass No
OpenML OpenML-ID-40668 Games 67557 43 3 multiclass No
OpenML OpenML-ID-40670 Other 3186 181 3 multiclass Yes
OpenML OpenML-ID-40701 Other 5000 21 2 binclass Yes
OpenML OpenML-ID-40923 Other 92000 1025 46 multiclass No
OpenML OpenML-ID-40927 Handwriting 60000 3073 10 multiclass No
OpenML OpenML-ID-40966 Other 1080 82 8 multiclass No
OpenML OpenML-ID-40975 Other 1728 7 4 multiclass Yes
OpenML OpenML-ID-40978 Other 3279 1559 2 binclass Yes
OpenML OpenML-ID-40979 Other 2000 241 10 multiclass Yes
OpenML OpenML-ID-40982 Other 1941 28 7 multiclass Yes
OpenML OpenML-ID-40983 Other 4839 6 2 binclass Yes
OpenML OpenML-ID-40984 Other 2310 20 7 multiclass Yes
OpenML OpenML-ID-40994 Other 540 21 2 binclass Yes
OpenML OpenML-ID-40996 Other 70000 785 10 multiclass No
OpenML OpenML-ID-41027 Games 44819 7 3 multiclass Yes
Table 14: Details of OpenML-CC18 benchmark datasets.
Table 15: TALENT benchmark datasets (auto-discovered, multiple domains).
Benchmark Dataset Name Domain Samples Features Classes Task Type Used In Experimentation
TALENT ASP-POTASSCO-class Other 1294 141 11 multiclass Yes
TALENT Amazon_employee_access Other 32769 7 2 binclass Yes
TALENT BLE_RSSI__Indoor_localization Other 9984 3 3 multiclass Yes
TALENT BNG(breast-w) Healthcare 39366 9 2 binclass Yes
TALENT BNG(cmc) Other 55296 9 3 multiclass Yes
TALENT BNG(tic-tac-toe) Other 39366 9 2 binclass Yes
TALENT Bank_Customer_Churn Finance 10000 10 2 binclass Yes
TALENT Basketball_c Retail 1340 11 2 binclass Yes
TALENT CDC_Diabetes_Health Healthcare 253680 21 2 binclass Yes
TALENT California-Housing-Class Other 20640 8 2 binclass No
TALENT Cardiovascular-Disease Healthcare 70000 11 2 binclass Yes
TALENT Click_prediction_small Other 39948 3 2 binclass Yes
TALENT Credit_c Finance 100000 22 3 multiclass Yes
TALENT Customer_Personality_Analysis Retail 2240 24 2 binclass Yes
TALENT DataScience_Kiva_Crowdfunding Other 671205 11 4 multiclass No
TALENT Diabetic_Retinopathy_Debrecen Healthcare 1151 19 2 binclass Yes
TALENT E-CommereShippingData Other 10999 10 2 binclass Yes
TALENT Employee Other 4653 8 2 binclass Yes
TALENT FICO-HELOC-cleaned Other 9871 23 2 binclass Yes
TALENT FOREX_audcad-day-High Finance 1834 10 2 binclass No
TALENT FOREX_audcad-hour-High Finance 43825 10 2 binclass No
TALENT FOREX_audchf-day-High Finance 1833 10 2 binclass No
TALENT FOREX_audjpy-day-High Finance 1832 10 2 binclass No
TALENT FOREX_audjpy-hour-High Finance 43825 10 2 binclass No
TALENT FOREX_audsgd-hour-High Finance 43825 10 2 binclass No
TALENT FOREX_audusd-hour-High Finance 43825 10 2 binclass No
TALENT FOREX_cadjpy-day-High Finance 1834 10 2 binclass No
TALENT FOREX_cadjpy-hour-High Finance 43825 10 2 binclass No
TALENT Firm-Teacher_Clave-Direction Other 10800 16 4 multiclass Yes
TALENT Fitness_Club_c Other 1500 6 2 binclass Yes
TALENT GAMETES_Epistasis_2-Way Games 1600 20 2 binclass Yes
TALENT GAMETES_Heterogeneity Games 1600 20 2 binclass Yes
TALENT Gender_Gap_in_Spanish Other 4746 13 3 multiclass Yes
TALENT GesturePhaseSegmentation Other 9873 32 5 multiclass Yes
TALENT HR_Analytics_Job_Change Other 19158 13 2 binclass Yes
TALENT IBM_HR_Analytics Other 1470 31 2 binclass Yes
TALENT INNHotelsGroup Other 36275 17 2 binclass Yes
TALENT Indian_pines Other 9144 220 8 multiclass Yes
TALENT JapaneseVowels Other 9961 14 9 multiclass Yes
TALENT KDDCup09_upselling Other 5128 49 2 binclass Yes
TALENT MIC Other 1649 104 2 binclass Yes
TALENT MagicTelescope Other 19020 9 2 binclass Yes
TALENT Marketing_Campaign Finance 2240 27 2 binclass Yes
TALENT Mobile_Price_Classification Telcom 2000 20 4 multiclass Yes
TALENT National_Health_and_Nutrition Healthcare 2278 7 2 binclass Yes
TALENT PhishingWebsites Other 11055 30 2 binclass Yes
TALENT PieChart3 Other 1077 37 2 binclass Yes
TALENT Pima_Indians_Diabetes Healthcare 768 8 2 binclass Yes
TALENT PizzaCutter3 Other 1043 37 2 binclass Yes
TALENT Pumpkin_Seeds Other 2500 12 2 binclass Yes
TALENT QSAR_biodegradation Healthcare 1054 41 2 binclass Yes
TALENT Rain_in_Australia Other 145460 18 3 multiclass No
TALENT SDSS17 Other 100000 12 3 multiclass Yes
TALENT Satellite Other 5100 36 2 binclass Yes
TALENT Smoking_and_Drinking Other 991346 23 2 binclass No
TALENT Telecom_Churn_Dataset Telcom $33$ 17 2 binclass Yes
TALENT UJI_Pen_Characters Other 1364 80 35 multiclass Yes
TALENT Water_Quality_and_Potability Manufacturing 3276 8 2 binclass Yes
TALENT Wilt Other 4821 5 2 binclass Yes
TALENT abalone Other 4177 8 3 multiclass Yes
TALENT accelerometer Other 153004 4 4 multiclass No
TALENT ada Other 4147 48 2 binclass Yes
TALENT ada_agnostic Other 4562 48 2 binclass Yes
TALENT ada_prior Other 4562 14 2 binclass Yes
TALENT airlines_seed_0_nrows Telcom 2000 7 2 binclass Yes
TALENT allbp Other 3772 29 3 multiclass Yes
TALENT allrep Other 3772 29 4 multiclass Yes
TALENT analcatdata_authorship Other 841 69 4 multiclass Yes
TALENT artificial-characters Other 10218 7 10 multiclass Yes
TALENT autoUniv-au4-2500 Other 2500 100 3 multiclass Yes
TALENT autoUniv-au7-1100 Other 1100 12 5 multiclass Yes
TALENT bank Finance 45211 16 2 binclass Yes
TALENT banknote_authentication Finance 1372 4 2 binclass Yes
TALENT baseball Other 1340 16 3 multiclass Yes
TALENT car-evaluation Other 1728 21 4 multiclass Yes
TALENT churn Finance 5000 20 2 binclass Yes
TALENT cmc Other 1473 9 3 multiclass Yes
TALENT company_bankruptcy_prediction Finance 6819 95 2 binclass Yes
TALENT compass Other 16644 17 2 binclass Yes
TALENT contraceptive_method_choice Other 1473 9 3 multiclass Yes
TALENT credit Finance 16714 10 2 binclass Yes
TALENT customer_satisfaction_in_airline Retail 129880 21 2 binclass No
TALENT dabetes_130-us_hospitals Healthcare 101766 20 2 binclass Yes
TALENT default_of_credit_card_clients Finance 30000 23 2 binclass Yes
TALENT delta_ailerons Other 7129 5 2 binclass Yes
TALENT dis Other 3772 29 2 binclass Yes
TALENT dna Healthcare 3186 180 3 multiclass Yes
TALENT drug_consumption Healthcare 1884 12 7 multiclass Yes
TALENT dry_bean_dataset Other 13611 16 7 multiclass Yes
TALENT eeg-eye-state Other 14980 14 2 binclass Yes
TALENT electricity Manufacturing 45312 8 2 binclass Yes
TALENT estimation_of_obesity_levels Other 2111 16 7 multiclass Yes
TALENT eye_movements Healthcare 10936 27 3 multiclass Yes
TALENT eye_movements_bin Healthcare 7608 20 2 binclass Yes
TALENT first-order-theorem-proving Other 6118 51 6 multiclass Yes
TALENT gas-drift Other 13910 128 6 multiclass Yes
TALENT gina_agnostic Other 3468 970 2 binclass Yes
TALENT golf_play_dataset_extended Other 1095 9 2 binclass Yes
TALENT heloc Other 10000 22 2 binclass Yes
TALENT hill-valley Other $12$ 100 2 binclass No
TALENT house_16H Other 13488 16 2 binclass Yes
TALENT htru Other 17898 8 2 binclass Yes
TALENT ibm-employee-performance Other 1470 30 2 binclass Yes
TALENT in_vehicle_coupon_recos Retail 12684 21 2 binclass Yes
TALENT internet_firewall Other 65532 7 4 multiclass Yes
TALENT internet_usage Other 10108 70 46 multiclass Yes
TALENT jm1 Other 10885 21 2 binclass No
TALENT jungle_chess_2pcs_raw Other 44819 6 3 multiclass Yes
TALENT kc1 Other 2109 21 2 binclass No
TALENT kdd_ipums_la_97-small Other 5188 20 2 binclass Yes
TALENT kr-vs-k Other 28056 6 18 multiclass Yes
TALENT kropt Other 28056 6 18 multiclass Yes
TALENT led24 Other 3200 24 10 multiclass Yes
TALENT led7 Other 3200 7 10 multiclass Yes
TALENT letter Other 20000 15 26 multiclass Yes
TALENT madeline Other 3140 259 2 binclass Yes
TALENT mammography Other 11183 6 2 binclass Yes
TALENT maternal_health_risk Healthcare 1014 6 3 multiclass Yes
TALENT mfeat-factors Other 2000 216 10 multiclass Yes
TALENT mfeat-fourier Other 2000 76 10 multiclass Yes
TALENT mfeat-karhunen Other 2000 64 10 multiclass Yes
TALENT mfeat-morphological Other 2000 6 10 multiclass Yes
TALENT mfeat-pixel Other 2000 240 10 multiclass Yes
TALENT mfeat-zernike Other 2000 47 10 multiclass Yes
TALENT mice_protein_expression Other 1080 75 8 multiclass Yes
TALENT microaggregation2 Other 20000 20 5 multiclass Yes
TALENT mobile_c36_oversampling Telcom 51760 6 2 binclass Yes
TALENT mozilla4 Other 15545 4 2 binclass Yes
TALENT naticusdroid+android Healthcare 29332 86 2 binclass Yes
TALENT national-longitudinal-survey-binary Other 4908 16 2 binclass Yes
TALENT okcupid_stem Other 26677 13 3 multiclass Yes
TALENT one-hundred-plants-margin Other 1600 64 100 multiclass Yes
TALENT one-hundred-plants-shape Other 1600 64 100 multiclass Yes
TALENT one-hundred-plants-texture Other 1599 64 100 multiclass Yes
TALENT online_shoppers Retail 12330 14 2 binclass No
TALENT optdigits Other 5620 64 10 multiclass Yes
TALENT ozone-level-8hr Other 2534 72 2 binclass Yes
TALENT page-blocks Other 5473 10 5 multiclass Yes
TALENT pc1 Other 1109 21 2 binclass No
TALENT pc3 Other 1563 37 2 binclass No
TALENT pc4 Other 1458 37 2 binclass No
TALENT pendigits Other 10992 16 10 multiclass Yes
TALENT phoneme Other 5404 5 2 binclass Yes
TALENT pol Other 10082 26 2 binclass Yes
TALENT predict_students_dropout Other 4424 34 3 multiclass Yes
TALENT qsar Other 1055 40 2 binclass Yes
TALENT rice_cammeo_and_osmancik Other 3810 7 2 binclass Yes
TALENT ringnorm Other 7400 20 2 binclass Yes
TALENT rl Other 4970 12 2 binclass No
TALENT satimage Other 6430 36 6 multiclass Yes
TALENT segment Other 2310 17 7 multiclass Yes
TALENT seismic+bumps Other 2584 18 2 binclass Yes
TALENT semeion Other 1593 256 10 multiclass No
TALENT shuttle Other 58000 9 7 multiclass Yes
TALENT spambase Other 4601 57 2 binclass Yes
TALENT splice Other 3190 60 3 multiclass Yes
TALENT sports_articles_for_objectivity Other 1000 59 2 binclass Yes
TALENT statlog Other 1000 20 2 binclass Yes
TALENT steel_plates_faults Other 1941 27 7 multiclass Yes
TALENT sylvine Other 5124 20 2 binclass Yes
TALENT taiwanese_bankruptcy Finance 6819 95 2 binclass Yes
TALENT telco-customer-churn Telcom 7043 18 2 binclass Yes
TALENT texture Other 5500 40 11 multiclass Yes
TALENT thyroid Healthcare 7200 21 3 multiclass Yes
TALENT thyroid-ann Healthcare 3772 21 3 multiclass Yes
TALENT thyroid-dis Healthcare 2800 26 5 multiclass Yes
TALENT turiye_student_evaluation Other 5820 32 5 multiclass Yes
TALENT twonorm Other 7400 20 2 binclass Yes
TALENT vehicle Other 846 18 4 multiclass Yes
TALENT volkert Other 58310 180 10 multiclass Yes
TALENT walking-activity Other 149332 4 22 multiclass No
TALENT wall-robot-navigation Other 5456 24 4 multiclass Yes
TALENT water_quality Manufacturing 7996 20 2 binclass Yes
TALENT waveform-5000 Other 5000 40 3 multiclass Yes
TALENT waveform_database_generator Other 4999 21 3 multiclass Yes
TALENT waveform_database_generator-v2 Other 5000 21 3 multiclass Yes
TALENT website_phishing Other 1353 9 3 multiclass Yes
TALENT wine Manufacturing 2554 4 2 binclass No
TALENT wine-quality-red Manufacturing 1599 4 6 multiclass Yes
TALENT wine-quality-white Manufacturing 4898 11 7 multiclass Yes
TALENT yeast Other 1484 8 10 multiclass Yes
Table 15: Details of TALENT benchmark datasets.
Table 16: TabZilla benchmark datasets (36 datasets via OpenML).
Benchmark Dataset Name Domain Samples Features Classes Task Type Used In Experimentation
TabZilla OpenML-ID-999 Health 226 70 2 binclass Yes
TabZilla OpenML-ID-10 Health 148 19 4 multiclass Yes
TabZilla OpenML-ID-11 Other 625 5 3 multiclass Yes
TabZilla OpenML-ID-14 Other 2000 77 10 multiclass Yes
TabZilla OpenML-ID-22 Other 2000 48 10 multiclass Yes
TabZilla OpenML-ID-29 Finance 690 16 2 binclass Yes
TabZilla OpenML-ID-27 Health 368 23 2 binclass Yes
TabZilla OpenML-ID-31 Finance 1000 21 2 binclass Yes
TabZilla OpenML-ID-46 Other 3190 61 3 multiclass Yes
TabZilla OpenML-ID-54 Other 846 19 4 multiclass Yes
TabZilla OpenML-ID-333 Other 556 7 2 binclass Yes
TabZilla OpenML-ID-1067 Other 2109 22 2 binclass Yes
TabZilla OpenML-ID-1468 Other 1080 857 9 multiclass Yes
TabZilla OpenML-ID-1494 Other 1055 42 2 binclass Yes
TabZilla OpenML-ID-43973 Other 3172 6 2 binclass Yes
TabZilla OpenML-ID-1043 Other 4562 49 2 binclass Yes
TabZilla OpenML-ID-43945 Other 38474 9 2 binclass Yes
TabZilla OpenML-ID-1486 Other 34465 119 2 binclass No
TabZilla OpenML-ID-42825 Other 8378 123 No
TabZilla OpenML-ID-4538 Other 9873 33 5 multiclass Yes
TabZilla OpenML-ID-23512 Other 98050 29 2 binclass No
TabZilla OpenML-ID-4134 Other 3751 1777 2 binclass Yes
TabZilla OpenML-ID-470 Other 672 10 2 binclass No
TabZilla OpenML-ID-1493 Other 1599 65 100 multiclass Yes
TabZilla OpenML-ID-1459 Other 10218 8 10 multiclass Yes
TabZilla OpenML-ID-41027 Games 44819 7 3 multiclass Yes
TabZilla OpenML-ID-40981 Other 690 15 2 binclass Yes
TabZilla OpenML-ID-934 Other 1156 6 2 binclass Yes
TabZilla OpenML-ID-1565 Health 294 14 5 multiclass Yes
TabZilla OpenML-ID-41150 Other 130064 51 2 binclass No
TabZilla OpenML-ID-41159 Other 20000 4297 2 binclass No
TabZilla OpenML-ID-846 Other 16599 19 2 binclass Yes
TabZilla OpenML-ID-1169 Other 539383 8 2 binclass No
TabZilla OpenML-ID-41147 Other 425240 79 2 binclass Yes
TabZilla OpenML-ID-41143 Other 2984 145 2 binclass Yes
TabZilla OpenML-ID-1567 Other 1025009 11 10 multiclass No
Table 16: Details of TabZilla benchmark datasets.
Note: Some TabZilla dataset statistics are to be completed. TabZilla datasets are accessed via OpenML and may include both dataset IDs and task IDs (TabZilla automatically handles task-to-dataset conversion).
Table 17: Datasets used for fairness evaluation (contain demographic/sensitive attributes).
Source Dataset Name Domain Samples Features Classes Sensitive Features
UCI Adult Census Income Other 48842 14 binclass Race, Sex, Age
OpenML German Credit (ID-31) Finance 1000 21 binclass Age, Personal-status-sex
Kaggle COMPAS Recidivism Legal  7000 53 binclass Race, Sex
UCI Default of Credit Card Clients (Taiwan) Finance 30000 23 binclass Sex, Education, Marital Status
UCI Student Performance in Exams (Portuguese) Education 649 33 binclass Gender, Race/Ethnicity, Parental Level of Education
Kaggle Law School Admission Legal  20000 12 binclass Race, Sex
Kaggle IBM HR Analytics Employee Attrition Other 1470 31 binclass Gender, Age, MaritalStatus, RelationshipSatisfaction
Offical HMDA API Home Mortgage Disclosure Act (NY 2023) Finance  32000 20+ multiclass Race ,Ethnicity, Sex, Age
Kaggle NHANES Demographics (2013-14) Health 5000 10+ multiclass Gender, Age, Race, Ethnicity
Table 17: Common datasets used for fairness evaluation.
Note: Fairness evaluation is performed on datasets where sensitive features are explicitly identified. The sensitive features column indicates demographic or protected attributes available in each dataset. Fairness metrics (SPD, EOD, EOpD) measure disparities in model predictions across different groups defined by these sensitive features.
12  Appendix C: Glossary of Key Terms
This glossary provides concise definitions of technical terms used throughout this paper. For detailed explanations, refer to the relevant sections indicated.
Table 18: Glossary of key technical terms.
Term Definition
Brier Score Metric quantifying both calibration and refinement by penalizing squared deviations between predicted probabilities and binary outcomes; lower is better (see Section 6 for details).
Equalized Odds Difference (EOD) Measure of disparities in both true positive and false positive rates across demographic groups; more stringent than statistical parity, requiring consistent error profiles (zero indicates perfect equity).
Equalized Opportunity Difference (EOpD) Focus on true positive rate parity across demographic groups; ensures qualified individuals receive equal treatment regardless of group membership (zero indicates perfect equity).
Episodic Training Meta-learning approach where each training step samples disjoint support and query sets from training data; mimics few-shot learning scenarios and preserves in-context generalization (see Section 2).
Expected Calibration Error (ECE) Average deviation between predicted confidence and observed accuracy across binned predictions; lower values indicate better calibration alignment (see Section 6).
Low-Rank Adaptation (LoRA) Parameter-efficient fine-tuning technique that adds trainable low-rank matrices to approximate full weight updates; only adapter matrices are updated, base weights remain frozen (see Section 2).
Maximum Calibration Error (MCE) Worst-case calibration gap across all prediction bins; identifies regions of systematic miscalibration and extreme confidence misalignment (see Section 6).
Meta-Learning (Episodic Fine-Tuning) Training strategy using support-query splits that mimics inference-time environment; preserves in-context generalization while enabling task-specific adaptation (see Section 2).
Parameter-Efficient Fine-Tuning (PEFT) Adaptation methods that update only a subset of model parameters, reducing computational costs by 60–80% while preserving performance; includes LoRA and similar techniques (see Section 2).
Statistical Parity Difference (SPD) Absolute difference in positive prediction rates across demographic groups; tests demographic parity, where zero indicates perfect parity in prediction distribution (see Section 6).
Supervised Fine-Tuning (SFT) Full parameter updates on labeled data using standard gradient descent; updates all model parameters, computationally intensive but can achieve highest task-specific accuracy (see Section 2).
Zero-Shot Inference Prediction using pretrained model weights without parameter updates; models receive training samples as context during inference, performing in-context learning through forward passes alone (see Section 2).